### v1

# SGD with Clipping is Secretly Estimating the Median Gradient

Abstract
There are several applications of stochastic optimization where one can benefit from a robust estimate of the gradient. For example, domains such as distributed learning with corrupted nodes, the presence of large outliers in the training data, learning under privacy constraints, or even heavy-tailed noise due to the dynamics of the algorithm itself. Here we study SGD with robust gradient estimators based on estimating the median. We first consider computing the median gradient across samples, and show that the resulting method can converge even under heavy-tailed, state-dependent noise. We then derive iterative methods based on the stochastic proximal point method for computing the geometric median and generalizations thereof. Finally we propose an algorithm estimating the median gradient across iterations, and find that several well known methods – in particular different forms of clipping – are particular cases of this framework.
Machine Learning, ICML
[…]
figuret
1 Introduction
Many problems in machine learning can be represented by
$\min_{\bm{w}\in\mathbb{R}^{d}}\ell(\bm{w}),$
an optimization problem where $\bm{w}$ is a vector of parameters and $\ell:\mathbb{R}^{d}\to\mathbb{R}$ is a loss function. To tackle this problem, gradient-based optimization algorithms have been the de facto choice in many application domains (Nemirovsky & Yudin, 1983; Bottou, 2010; Bottou et al., 2018), where we are only given access to a noisy oracle of the gradient, i.e.,
$\bm{g}_{t}=\nabla\ell(\bm{w}_{t})+\bm{\xi}_{t}(\bm{w}_{t}).$
Here $t$ denotes the iterations and $\bm{\xi}_{t}(\bm{w}_{t})\in\mathbb{R}^{d}$ is noise vector that is sampled from a distribution that depends on the parameter $\bm{w}_{t}$. With this noise oracle, stochastic gradient descent (SGD) with learning rate $\eta_{t}$ is given by
$\bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\bm{g}_{t}.$ (SGD)
While SGD has proven useful in numerous applications, its performance heavily relies on the ‘quality’ of the noisy oracles $\bm{g}_{t}$, and can tragically degrade when the noise $\bm{\xi}_{t}$ has significant outliers, exhibits heavy tails (i.e., $\|\bm{\xi}_{t}\|$ can be very large with non-negligible probability), or has been corrupted in an adversarial way. We provide two examples.
  1. (i)
    Heavy-tailed gradients. Recent studies have provided theoretical and empirical evidence that heavy tails can naturally arise in stochastic optimization. From a theoretical perspective (Gurbuzbalaban et al., 2021; Hodgkinson & Mahoney, 2021; Pavasovic et al., 2023) showed that the parameter-dependent nature of $\bm{\xi}_{t}$ can result in heavy tails, even in very basic settings such as linear regression with Gaussian data. On the other hand, it has been empirically observed that the gradients for training transformer architectures on language tasks are more heavy-tailed compared to, for example, convolutional models for image data (Zhang et al., 2020b; Kunstner et al., 2023). While the presence of heavy-tails in gradients might be beneficial in certain settings (Simsekli et al., 2020), from an optimization perspective, it mostly introduces non-trivial challenges, which might even make the algorithm diverge unless additional care is taken (Zhang et al., 2020b; Gorbunov et al., 2020).
  2. (ii)
    Corrupted nodes. A well-known problem in distributed learning is when some computation nodes are malicious and can communicate adversarial updates, which results in inaccurate gradient oracles that can significantly misguide the optimization procedure. Similar to the heavy-tailed setting, depending on how much malicious are the nodes, SGD can be impractical unless certain modifications are made (Mhamdi et al., 2018). We refer to (Karimireddy et al., 2021) and references therein for an overview of more robust aggregation techniques.
In order to make the optimization algorithms more robust to such noisy oracles, many techniques have been developed for different domains under various conditions. In the context of heavy-tailed gradients, clipping is often employed (Zhang et al., 2020a; Puchkin et al., 2023; Koloskova et al., 2023). In distributed learning robust aggregation rules are needed to protect against corrupted or malicious nodes. (Data et al., 2018; Karimireddy et al., 2021; Khirirat et al., 2023). This line of research is also closely related to error feedback in federated learning (Seide et al., 2014; Karimireddy et al., 2019; Richtárik et al., 2021) and learning under differential privacy constraints (Khirirat et al., 2023).
Robustness of median.
While attracting increasing attention in stochastic optimization thanks to modern machine learning applications, taming the impact of strong noise has already been considered in various other domains. Indeed, estimation under heavy-tails and corruptions has been a long-standing topic in robust statistics (Huber, 1981). Being a vast research field that has produced numerous algorithms designed for different tasks, one of the recurring themes in robust statistics is the use of the sample median and its variations as opposed to the sample mean (Minsker, 2015), whenever an aggregation of random variables is needed. For instance, in the presence of heavy tails, the sample median has been shown to be a significantly more robust estimator of the true mean, whereas the sample mean can be vulnerable to strong noise (Lugosi & Mendelson, 2019).
Sample median SGD.
Inspired by tools from robust statistics, some notions of median have been utilized in stochastic optimization as well (Acharya et al., 2022; Yin et al., 2018). One such approach is to compute a sample median over a sample of gradients. That is, let
$\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\xi}^{(i)}_{t}(\bm{w}_{t})\quad\mbox{for}\quad i=1,\dots,n,$
where $\{\bm{\xi}^{(i)}_{t}(\bm{w}_{t})\in\mathbb{R}^{d}\}_{i=1}^{n}$ is a set of i.i.d. noise vectors.
Given these sampled gradients, stochastic optimization algorithms typically aggregate them by computing their sample mean, or average $\bm{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\bm{g}_{t}^{(i)}$, and then use this $\bm{g}_{t}$ in (SGD). Alternatively, to improve robustness, one can aggregate these gradients $\{\bm{g}_{t}^{(i)}\}_{i=1}^{n}$ using a multidimensional variant of the median. While this approach has paved the way for powerful optimization algorithms in terms of robustness, it often introduces a significant computational burden, since a multivariate median needs to be computed at every iteration (Weiszfeld & Plastria, 2009; Vardi & Zhang, 2000).
Motivation. Even though the clipping-based approaches developed in optimization theory and the median-based approaches inspired from robust statistics aim at solving similar tasks, they where designed in different mathematical frameworks under different conditions. In this study, our main goal is to bridge the gap between these two seemingly different branches and to bring a unified theoretical perspective that can shed further light on both directions by leveraging tools from both optimization and robust statistics.
Contributions. We first revisit median gradient-based optimization and consider a generic notion of median which include several multivariate medians that have been extensively used in practice. By exploiting the statistical properties of the sample median estimator, we show that SGD with (approximately computed) median gradients indeed yields a very powerful algorithm, which converges with $1/\sqrt{T}$ rate ($T$ being the number of iterations) even when the gradient oracles are heavy-tailed with diverging second-order moments. We further illustrate that, in addition to having infinite variance, the noise vectors $\bm{\xi}_{t}(\bm{w})$ can even have a multiplicative dependence on $\bm{w}$, a scenario which is highly challenging and cannot be directly covered by existing theoretical results, see e.g., (Wang et al., 2021; Zhang et al., 2020b; Puchkin et al., 2023).
Encouraged by this result, we then focus on the practical computation of the medians. To this end, we solve the optimization-based formulation of the median using Stochastic Proximal Point (SPP) (Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019; Toulis et al., 2020). As an alternative, we also consider iterative solvers based on the stochastic subgradient method.
We then depart from previous work by approximately computing the median gradient across iterations. That is, the median over $\{\bm{g}_{t}\}_{t=0}^{T}.$ We use this online median gradient within the SGD framework. When using the Stochastic Proximal Point method for computing the median across iterations, we recover several known clipping-based optimization algorithms as special cases. In particular, we recover different forms of clipping (Gorbunov et al., 2020; Zhang et al., 2020a), such as Clip21 for distributed learning (Khirirat et al., 2023), and signed gradient methods (Riedmiller & Braun, 1993; Bernstein et al., 2018). We also show that (non-clipped) heavy-ball momentum (Polyak, 1964) is in fact an online mean estimator.
Our proposed framework allows to unify the different motivations for previously developed techniques, and illustrates that clipping-based optimization algorithms essentially runs a “hidden” estimation of the median across iterations. In particular, the observation that centered clipping (Karimireddy et al., 2021) and Clip21 are an iterative estimator of a geometric median appears to be new, and might have consequences for its application in distributed learning.
We finally illustrate our theory on experiments where we compare the effectiveness of sample median, sample mean, and several online median estimators. Our results underline that for heavy-tailed noise, using the sample median is highly effective in contrast to the sample mean which is unstable and often does not converge. Our experiments also show that our online median estimates are robust, while being less expensive to compute. We further compare different clipping techniques for training transformer architectures on language modeling tasks, and show that they improve upon the performance of SGD with momentum.
[…]
2 Preliminaries
2.1 Heavy-tailed Random Variables
A random variable $X$ is said to be heavy-tailed if the tails of its distribution decay slower than the tails of an exponential distribution. One important example, is the family of symmetric $\alpha$-stable distributions (Nolan, 2020, Def. 1.4), denoted $\mathcal{P}_{\alpha,\sigma}$: it is parameterized by a stability parameter $\alpha\in(0,2]$ and a scale parameter $\sigma>0$. For $\alpha=1$, it is the Cauchy distribution, while for $\alpha=2$ it is a Gaussian. The parameter $\sigma$ controls the spread of the distribution. When $\sigma=1$, we call the distribution standard and denote it by $\mathcal{P}_{\alpha}:=\mathcal{P}_{\alpha,1}$. The parameter $\alpha$ on the other hand controls the heaviness of the tails: for $\alpha<2$ and $X\sim\mathcal{P}_{\alpha}$ has infinite variance, whereas for $\alpha\leq 1$ even the mean of $X$ is infinite. However, for any $\alpha$, the median is equal to 00.
2.2 Mean Estimation and (Sample) Median
Estimating the mean of a distribution, given a finite number of samples, is one of the core problems studied in statistics. When the distribution in question has heavy-tails, or has significant outliers, the sample mean turns out to be a poor estimator. We refer to (Lugosi & Mendelson, 2019) for a survey of alternative techniques. In this section, we will introduce basic definitions for median.
In this section we denote by $z$ a real-value random value, and with $\bm{z}\in\mathbb{R}^{d}$ a random vector. We always denote the expectation with respect to the distribution of $z$ (or $\bm{z}$ respectively) by $\mathbb{E}$.
Multivariate median.
The one-dimensional median is defined by
$\texttt{median}{}(z)\in\underset{m\in\mathbb{R}}{\operatorname{argmin}}\;\mathbb{E}\left[|m-z|\right].$ (1)
In contrast to the mean, the median is always defined (though it may not be unique) (Cramér, 2016).
Since we are interested in approximating the median gradient, we need a notion of median that generalizes to random vectors $\bm{z}\in\mathbb{R}^{d}$. For this, we can define the $\ell_{p}$-median:
$\texttt{median}_{\ell_{p}}(\bm{z})\in\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathbb{E}\left[\|\bm{m}-\bm{z}\|_{p}\right].$ (2)
This recovers several existing notions of a multivariate median: for $p=1$ we obtain a componentwise median (or $\ell_{1}$-median) and for $p=2$ we obtain the geometric median (or $\ell_{2}$-median) (Fréchet, 1948; Weiszfeld & Plastria, 2009; Minsker, 2015; Cohen et al., 2016). Clearly, for $d=1$ all notions conincide with the definition in (1).
Sample median.
Given $n$ realizations $\bm{z}^{(1)},\ldots,\bm{z}^{(n)}$ of $\bm{z}$, the sample median is defined as
$\texttt{median}_{\ell_{p},i\in[n]}(\bm{z}^{(i)})\in\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\frac{1}{n}\sum_{i=1}^{n}\|\bm{m}-\bm{z}^{(i)}\|_{p}.$ (3)
We remark that (3) is a special case of (2), by setting the distribution of $\bm{z}$ as a sum of Dirac measures $\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{z}_{i}}$.
It is generally accepted that the sample median is a more robust estimator than the sample mean (Huber, 1981; Hampel et al., 2005). This is nicely illustrated by the notion of a breakdown point (Donoho & Huber, 1983; Lopuhaa & Rousseeuw, 1991), which is the smallest fraction of a sample that, if arbitrarily corrupted, can arbitrarily change the value of the estimator. The breakdown point of the sample median is roughly $\tfrac{1}{2}$, meaning that at least half of the samples need to be outliers for the median to diverge. The sample mean however can be arbitrarily corrupted by a single outlier, thus its breakdown point is one over $\tfrac{1}{n}$.
If the samples $\bm{z}^{(i)}$ are obtained by taking block-wise means over observations, (3) is called the median-of-means estimator (Nemirovsky & Yudin, 1983; Lugosi & Mendelson, 2019), which is known to be a robust mean estimator (Lugosi & Mendelson, 2019).
Relation to M-estimation.
More generally, for a function $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ we will later consider the problem
$\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right].$ (4)
We assume that $\mathcal{D}$ is a closed, proper, convex function and that (4) admits a solution (see Appendix B for a discussion). For $\mathcal{D}=\tfrac{1}{2}\|\cdot\|_{2}^{2}$, the solution to (4) is clearly the mean $\mathbb{E}\left[\bm{z}\right]$, and for $\mathcal{D}=\|\cdot\|_{2}$, the solution is the geometric median as defined in (2). Problem (4) can also be seen as a special case of M-estimators (Huber, 1981, Sec. 3.2).
The choice of $\mathcal{D}$ determines properties of the associated estimator: if the function $\mathcal{D}$ heavily penalizes large values, like the quadratic $\mathcal{D}=\tfrac{1}{2}\|\cdot\|_{2}^{2}$, then the solution to (4) will be sensitive to outliers. In contrast $\mathcal{D}=\|\cdot\|_{2}$ and $\mathcal{D}=\|\cdot\|_{1}$ grow only linearly, and thus (4) will be less sensitive to outliers. We return to problem (4) later in Section 4.
For SGD, its update direction is a sampled gradient $\bm{g}_{t}$ which aims to approximate the full gradient $\nabla\ell(\bm{w}_{t})$. This section motivates that in the presence of large outliers it might be necessary to use more robust gradient estimators as update direction. A natural alternative is to use the sample median together with SGD (Alistarh et al., 2018; Acharya et al., 2022), as we do next.
3 Warm-up: Sample Median Gradient
In this section, we showcase that SGD with an (approximate) sample median is robust to heavy-tailed noise, whereas SGD using the sample mean can fail. Consider the following setup: at every iteration $t\geq 0$, we have access to $n$ noisy oracle gradients $\big{(}\bm{g}_{t}^{(1)},\ldots,\bm{g}_{t}^{(n)}\big{)}$, and we apply
$\bm{w}_{t+1}=\bm{w}_{t}-\eta(\bm{m}_{t}+\bm{e}_{t}),$ (5)
where $\bm{m}_{t}=\texttt{median}_{\ell_{p},i\in[n]}(\bm{g}_{t}^{(i)})$, and $\bm{e}_{t}\in\mathbb{R}^{d}$ represents an error we may commit in computing the sample median.
Assumption 3.1.
There exists $\sigma_{1},\sigma_{2},\delta_{1},\delta_{2}\geq 0$ for every $\bm{w}\in\mathbb{R}^{d}$, such that the sample median $\bm{m}=\texttt{median}_{\ell_{p},i\in[n]}(\bm{g}^{(i)})$ verifies
$\begin{split}&\|\mathbb{E}\left[\bm{m}\right]-\nabla\ell(\bm{w})\|^{2}\leq\delta_{1}^{2}+\delta_{2}\|\nabla\ell(\bm{w})\|^{2},\\ &\mathbb{E}\left[\|\bm{m}-\mathbb{E}\left[\bm{m}\right]\|^{2}\right]\leq\sigma_{1}^{2}+\sigma_{2}\|\nabla\ell(\bm{w})\|^{2}.\end{split}$ (6)
Assumption 3.2.
There exists $\varepsilon>0$ such that, for every $t\geq 0$, the error in computing the sample median has finite variance: $\mathbb{E}\left[\|\bm{e}_{t}\|^{2}\ |\ \bm{g}_{t}^{(1)},\dots,\bm{g}_{t}^{(n)}\right]\leq\varepsilon^{2}$.
Based on these two assumptions, we establish the following complexity results for (5).
Proposition 3.3.
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be $L$-smooth and let Assumption  3.1 hold with $\delta_{2}\leq\frac{1}{8}$. Take $\eta\leq\frac{1}{8L(1+2\sigma_{2}+2\delta_{2})}$ and consider $\bm{w}_{t}$ a sequence generated by (5) where $\bm{e}_{t}$ verifies Assumption 3.2. Then, for every $T\geq 1$ it holds
$\min_{0\leq t\leq T-1}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]=\mathcal{O}\left(\tfrac{1}{\eta T}+\eta+\delta_{1}^{2}+\varepsilon^{2}\right).$
Proposition 3.4.
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $G$-Lipschitz, with ${\rm{argmin}}~{}\ell\neq\emptyset$. Let Assumptions 3.1 and Assumption 3.2 hold. Without loss of generality assume that $\delta_{2}=\sigma_{2}=0$. Then, for $\eta=\frac{1}{(\delta+\varepsilon)T+\sqrt{T}}$ and every $T\geq 1,$ the iterates $\bm{w}_{t}$ from (5) satisfy
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]=\mathcal{O}\left(\frac{1}{\sqrt{T}}+\delta_{1}+\varepsilon\right),$
where $\bar{\bm{w}}_{T}:=\tfrac{1}{T}\sum_{t=0}^{T-1}\theta^{t}\bm{w}_{t}$ with $\theta=\tfrac{T}{T+2}$.
The proofs are provided in Appendix A. Let us now discuss our main Assumption 3.1. Assume for simplicity that $\delta_{2}=\sigma_{2}=0$, that $\bm{e}_{t}=0$, and consider that $\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\xi}_{t}^{(i)}$, where $\bm{\xi}_{t}^{(i)}$ is a noise term. Then, Assumption 3.1 becomes [^1] See Lemma A.2 in the appendix for a proof.
$\mathbb{E}\left[\|\texttt{median}_{\ell_{p},i\in[n]}(\bm{\xi}_{t}^{(i)})\|^{2}\right]\leq\nu^{2}<+\infty.$ (7)
In other words, we need the second moment of the sample median of $\bm{\xi}_{t}^{(i)}$ to be uniformly bounded. We want to stress that our choice of using the median is crucial when the $\bm{\xi}_{t}^{(i)}$ come from a heavy-tailed distribution, such as a (univariate) standard $\alpha$-stable distribution, with $\alpha\in(1,2)$. Indeed, in this case we know that the variance of the sample median is finite for sufficiently large sample size $n$ (see (Bickel, 1967, Theorem 2.2) and (Nolan, 2020, Theorem 1.2)) [^2] We know more: The sample median of a parent distribution with median $\xi$ is asymptotically normal with location $\xi$ (Chu & Hotelling, 1955). For Cauchy distributions ($\alpha=1$), the sample median has a finite variance for $n\geq 5$ (Chu & Hotelling, 1955, Remark 1), while first and second moment of the sample mean are undefined.. On the contrary, if we were taking $\bm{m}_{t}$ as a more classical sample mean we would be in trouble, because the variance of the sample mean would be infinite [^3] This is due to the fact that, by definition of $\alpha$-stability, the sample mean of i.i.d sample follows the same $\alpha$-stable distribution (up to affine translation). everywhere. See Fig. 1 for an illustration of this phenomenon.
This simple discussion illustrates that Assumption 3.1 can be satisfied under mild conditions on the noise. We exploit this in the next corollary, which shows that the $\ell_{1}$-sample median can even accommodate state-dependent noise. Let us define first our model for such noise.
Assumption 3.5.
For every $\bm{w}\in\mathbb{R}^{d}$, we have that $\bm{g}=\nabla\ell(\bm{w})+\bm{\Sigma}(\bm{w})\bm{\zeta}$, where $\bm{\Sigma}(\bm{w})\in\mathbb{R}^{d\times d}$ is a matrix and $\bm{\zeta}$ is a random vector whose components are drawn i.i.d. from a standard symmetric $\alpha$-stable distribution with $\alpha~{}>~{}1$. There exists constants $c_{1},c_{2}>0$ such that
$\|\bm{\Sigma}(\bm{w})\|_{F}^{2}\leq c_{1}+c_{2}\|\nabla\ell(\bm{w})\|^{2}\quad\mbox{for all }\bm{w}\in\mathbb{R}^{d}.$
As opposed to the following assumption that has been commonly made in the recent literature (see e.g., (Zhang et al., 2020b; Puchkin et al., 2023; Sadiev et al., 2023)) $\mathbb{E}\left[\|\bm{g}-\nabla\ell(\bm{w})\|^{p}\right]\leq\sigma^{2}$ for some $p<2$ where $\sigma^{2}>0$ cannot depend on $\bm{w}$, Assumption 3.5 can accommodate much stronger noise. To illustrate this, consider $d=1$ and $\ell(\bm{w})=\bm{w}^{2}/2$. Assumption 3.5 allows the noise model $\bm{\Sigma}(\bm{w})\bm{\zeta}$ with $\bm{\Sigma}(\bm{w})=\sqrt{c_{1}+c_{2}\bm{w}^{2}}$, which results in strong noise, since $\bm{\zeta}$ has infinite variance if $\alpha<2$, and this is amplified by $\bm{\Sigma}(\bm{w})$ when $\bm{w}$ is large. Such strong noise makes vanilla SGD diverge, as we will illustrate in our experiments.
Corollary 3.6.
Let $n\geq 3$ be odd and $p=1$. Consider the iterates from (5) with $\bm{e}_{t}=0$, and suppose that $\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\Sigma}(\bm{w}_{t})\bm{\zeta}_{t}^{(i)}$ verifies Assumption 3.5.
  1. (i)
    If $\ell$ is smooth and $\eta=1/\sqrt{T}$ then we have that $\min_{0\leq t\leq T-1}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]=\mathcal{O}(1/\sqrt{T})$.
  2. (ii)
    If $\ell$ is convex and Lipschitz and $\eta=1/\sqrt{T}$ then we have that $\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]=\mathcal{O}(1/\sqrt{T})$, where $\bar{\bm{w}}_{T}$ is defined as in Proposition 3.4.
This result illustrates how powerful the sample median can be when used in gradient-based optimization. However, the method (5) has clear drawbacks: (i) it requires to draw multiple samples per iteration (multiple calls of the oracle), and (ii) we need to compute the sample median up to some tolerance. For instance, the geometric median can not be computed in closed form and there exists a vast body of literature on how to approximate it (Cohen et al., 2016).
We tackle (ii) in the next section by developing an efficient and inexpensive iterative method for approximately solving the general problem (4) which includes the sample median as a special case. We tackle issue (i) later in Section 5.
Figure 1: First and second moment of sample median/mean for standard ($\alpha=1.1$)-stable distribution. We approximate $\mathbb{E}$ by running $1000$ trials. Even for small sample size, the sample median has small variance (cf. (7)), in contrast to the sample mean.
4 Iterative Methods for Median Estimation
We now focus on the following gradient estimation problem: let $\mathcal{G}$ denote the (fixed) distribution over noisy gradients $\bm{g}$, and consider the problem of approximating
$\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\mathcal{D}(\bm{m}-\bm{g})\right].$ (8)
In the previous section, we used the sample median. It follows that solving (8) with $\mathcal{D}=\|\cdot\|_{p}$, and with $\mathcal{G}$ being the empirical distribution over a sample $(\bm{g}^{(1)},\ldots,\bm{g}^{(n)})$, the solution to (8) is the sample median (3).
For solving (8), we consider only stochastic iterative methods, that can be applied for any distribution $\mathcal{G}$. The most commonly used such method is SGD, which has been used for computing the median (Cardot et al., 2011). But here we advocate for a different method: The stochastic proximal point method (SPP(Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019). In cases where SPP has a closed form update, it is often preferred to SGD since is easier to tune (Asi & Duchi, 2019; Milzarek et al., 2022). Fortunately, as we show next, the SPP method enjoys closed-form updates when $\mathcal{D}\in\{\|\cdot\|_{1},\|\cdot\|_{2},\tfrac{1}{2}\|\cdot\|^{2}\}$.
The update of SPP with stepsize $\tau>0$ applied to (8) can be written as (see details in Appendix B.3)
$\bm{m}_{t+1}\;=\;\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t}),$ (9)
where we denote by $\mathrm{prox}_{\mathcal{D}}$ the proximal operator of $\mathcal{D}$,
$\mathrm{prox}_{\mathcal{D}}(\bm{m}):=\underset{\bm{y}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathcal{D}(\bm{y})+\frac{1}{2}\|\bm{y}-\bm{m}\|^{2}.$
Because we are interested in the $\ell_{p}$-median, we first characterize the SPP update (9) for when $\mathcal{D}$ is the $\ell_{p}$-norm, and show it simply require to project the sampled gradient onto a certain ball of radius $\tau$ centered at $\bm{m}_{t}$.
Proposition 4.1.
Let $\mathcal{D}=\|\cdot\|_{p}$ be the $\ell_{p}$-norm for $p\in[1,\infty].$ Let $q\in[1,\infty]$ such that $\frac{1}{p}+\frac{1}{q}=1$, and let $\mathbb{B}_{q}(\bm{m},\tau)\;:=\;\left\{\bm{y}\;:\;\|\bm{m}-\bm{y}\|_{q}\leq\tau\right\}.$ The SPP update (9) is given by
$\bm{m}_{t+1}$ $=\bm{m}_{t}\;+\;\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0},\tau)}\big{(}\bm{g}_{t}-\bm{m}_{t}\big{)}$ (10)
$=\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m}_{t},\tau)}\big{(}\bm{g}_{t}\big{)}.$ (11)
Figure 2: One SPP update (11) for $\mathcal{D}=\|\cdot\|_{1}$ (blue) or $\mathcal{D}=\|\cdot\|_{2}$ (green) amounts to project onto a ball of the dual norm, centered at the current iterate $\bm{m}_{t}$ and of radius $\tau>0$. The thick lines represent the projections of $\bm{g}_{t}$ for all possible values of $\tau$.
Using Proposition 4.1 we can now develop closed form updates of the SPP method for when $\mathcal{D}$ is the $\ell_{1}$- or the $\ell_{2}$-norm, with both methods being related to clipping.
Vectorwise clipping.
If we choose $\mathcal{D}=\|\cdot\|_{2}$, SPP performs projections onto the $\ell^{2}$-ball which is also known as clipping (Khirirat et al., 2023; Koloskova et al., 2023).
Corollary 4.2.
For $\mathcal{D}=\|\cdot\|_{2}$ update (10) is given by
$\bm{m}_{t+1}$ $=\bm{m}_{t}+\mathrm{clip}_{\tau,2}(\bm{g}_{t}-\bm{m}_{t}),$ (12)
where $\mathrm{clip}_{\tau,2}(\bm{v}):=\frac{\tau}{\max\{\tau,\|\bm{v}\|_{2}\}}\bm{v}$.
This corollary shows that the SPP method for computing the $\ell_{2}$-median can be written as $\bm{m}_{t+1}=\beta_{t}\bm{m}_{t}+(1-\beta_{t})\bm{g}_{t}$, with $\beta_{t}:=1-\tfrac{\tau}{\max\{\tau,\|\bm{g}_{t}-\bm{m}_{t}\|_{2}\}}$. In this sense, (12) is momentum where the momentum coefficient depends on how far $\bm{g}_{t}$ is from the previous estimate $\bm{m}_{t}$. This allows the iterates to be more robust to outliers, as samples which are too far from $\bm{m}_{t}$ will be down weighted (see also Fig. 2).
Componentwise clipping.
If we choose $\mathcal{D}=\|\cdot\|_{1}$, SPP performs projections onto the $\ell^{\infty}$-ball, also known as componentwise clipping. Here as well, clipping protects against large individual entries by shrinking them independently.
Corollary 4.3.
For $\mathcal{D}=\|\cdot\|_{1}$ update  (10) is given by
$\bm{m}_{t+1}=\bm{m}_{t}+\mathrm{clip}_{\tau,1}(\bm{g}_{t}-\bm{m}_{t}),$ (13)
where $\mathrm{clip}_{\tau,1}(\bm{v}):=\left(\min\{\max\{\bm{v}_{i},-\tau\},\tau\}\right)_{i=1}^{d}$.
SPP and momentum.
It is easy to see that the solution to problem (8) with $\mathcal{D}=\tfrac{1}{2}\|\cdot\|^{2}$ is the mean $\mathbb{E}\left[\bm{g}\right]$. Computing (9) for this choice of $\mathcal{D}$, we obtain
$\bm{m}_{t+1}=(1-\tfrac{\tau}{1+\tau})\bm{m}_{t}+\tfrac{\tau}{1+\tau}\bm{g}_{t}.$ (14)
Interestingly, this is the (heavy-ball) momentum gradient estimator (Polyak, 1964; Sutskever et al., 2013), where the momentum coefficient is given by $\beta=1/(1+\tau)\in[0,1).$
Rates for SPP
Now that we have presented how to compute the iterates (9) in practice, let us state its convergence properties. Since all these methods are instantiates of the SPP method, we can use standard results such as (Davis & Drusvyatskiy, 2019, Thm. 4.4).
Proposition 4.4.
Let $\mathcal{D}$ be a norm, let $\varphi(\bm{m}):=\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\mathcal{D}(\bm{m}-\bm{g})\right]$, and let $\bm{m}_{*}\in\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\varphi(\bm{m})$. If $\bm{m}_{t}$ is generated by (9) then for $\bar{\bm{m}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{m}_{t}$ we have
$\mathbb{E}\left[\varphi(\bar{\bm{m}}_{T})-\varphi(\bm{m}_{*})\right]=O\left(\frac{\|\bm{m}_{0}-\bm{m}_{*}\|^{2}}{\tau T}+\tau\right).$
In Appendix B we give further examples for $\mathcal{D}$, such as the Huber loss. We also compare SPP with SGD for solving (8), showing that sign-SGD is related to $\ell_{1}$-median estimation.
5 Gradient Estimator in the Wild
Our final objective is to use robust gradient estimators within a SGD-type method. Because at each iteration the weights $\bm{w}_{t}$ are updated, the distribution of the gradients (denoted by $\mathcal{G}$ in the previous section) also changes at each iteration of SGD. Our strategy is to interweave updates in the weights $\bm{w}_{t}$ with SPP updates in the gradient estimators $\bm{m}_{t}.$ That is, we sample one stochastic gradient $\bm{g}_{t}$ per iteration and set
$\begin{array}[]{llr}\bm{m}_{t+1}&=\quad\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t}),\\ \bm{w}_{t+1}&=\quad\bm{w}_{t}-\eta_{t}\bm{m}_{t+1}.\\ \end{array}$ (17)
Formally, $\bm{m}_{t}$ is now an estimate of the median of the gradients across iterations, which is in contrast to (5), where $\bm{m}_{t}$ was an approximation of the sample median at fixed $\bm{w}_{t}$. With the results from Section 4, we immediately obtain the following special cases (here $p\in\{1,2\}$):
(i)  if $\mathcal{D}=\frac{1}{2}\|\cdot\|^{2}$ :  $\begin{cases}\bm{m}_{t+1}=(1-\tfrac{\tau}{1+\tau})\bm{m}_{t}+\tfrac{\tau}{1+\tau}\bm{g}_{t}\\ \bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\bm{m}_{t+1},\end{cases}$
(ii)  if $\mathcal{D}=\|\cdot\|_{p}$ :  $\begin{cases}\bm{m}_{t+1}=\bm{m}_{t}+{\rm clip}_{p,\tau}(\bm{g}_{t}-\bm{m}_{t})\\ \bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\bm{m}_{t+1}.\end{cases}$
Equation (i) is SGD with momentum (14); notice how the SPP step size $\tau$ defines the momentum coefficient. Instead, (ii) corresponds to clipping, and here the SPP step size determines the clipping threshold. In other words, SGD with momentum is to online mean estimation what clipping is to online median estimation.
A slightly different approach would be to restart $\bm{m}_{t}=\bm{0}$ in every iteration (cold start). In this case, (i) reduces to SGD with no momentum, while in (ii) the clipping operator is applied to $\bm{g}_{t}$, instead of $\bm{g}_{t}-\bm{m}_{t}$, which is also a common technique in practice (Zhang et al., 2020a). We discuss this relation below in more detail.
Connection to Error Feedback.
Methods such as (17) are also used in the distributed learning setting to improve communication and protect against corrupted nodes (Karimireddy et al., 2021). These methods are often called Error Feedback (EF) methods (Seide et al., 2014; Richtárik et al., 2021) and generally involve an update of the form $\bm{m}_{t+1}=\bm{m}_{t}+\mathcal{C}_{t}(\bm{g}_{t}-\bm{m}_{t})$ where $\mathcal{C}_{t}:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ is a compression operator. Projections onto balls can be seen as a compression, since one can encode the resulting vector in fewer bits (Safaryan et al., 2021). In particular, centered clipping from (Karimireddy et al., 2021), and the Clip21 method (Khirirat et al., 2023) are using update (12) in the distributed setting, and are thus secretly estimating the geometric median of the gradients returned across all nodes.
SGD with standard clipping.
Even in the non-distributed setting, a standard technique to avoid training instabilities is to directly clip the gradient $\bm{g}_{t}$ (Zhang et al., 2020a). For momentum coefficient $\beta\in[0,1)$ and $c>0$, let
$\begin{split}\bm{m}_{t+1}&=\beta\bm{m}_{t}+(1-\beta)\min\{1,\tfrac{c}{\|\bm{g}_{t}\|}\}\bm{g}_{t},\\ \bm{w}_{t+1}&=\bm{w}_{t}-\eta_{t}\bm{m}_{t+1}.\end{split}$ (18)
We will refer to (18) as clipped-SGD. The method has been studied extensively, often for the case $\beta=0$, see (Zhang et al., 2020a; Gorbunov et al., 2020; Koloskova et al., 2023; Puchkin et al., 2023) and references therein.
We present two motivations for using clipped-SGD. If $\beta=0$ then (18) is the same as (12) with cold start (as explained above). When $\beta\neq 0$, clipped-SGD can also be viewed as a SPP method for computing a robust mean estimator given by
$\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\tfrac{1}{2}\|\bm{m}-\min\{1,\tfrac{c}{\|\bm{g}\|}\}\bm{g}\|^{2}\right].$ (19)
The solution to (19) is $\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\min\{1,\tfrac{c}{\|\bm{g}\|}\}\bm{g}\right]$, which is known as the Catoni-Giulini mean estimator (Catoni & Giulini, 2018; Lugosi & Mendelson, 2019). One step of SPP (9) applied to (19) gives (18) with $\beta=\tfrac{1}{1+\tau}$.
6 Experiments
We compare the iterative gradient estimation techniques for two experimental setups, namely (i) estimating a vector/gradient from a fixed distribution, and (ii) methods of the form (17), that is, iterative gradient estimation while simultaneously learning the weights.
We study both settings for synthetic data where we can control the level of heavy-tailedness, and for language modelling tasks with transformer models. In the following, we use VClip to refer to vectorwise clipping (12) and CClip for componentwise clipping  (13).
We consider the same three language modeling tasks as studied in (Kunstner et al., 2023): an encoder-only transformer for the PTB, a Transformer-XL model for the WikiText-2, and fine-tuning a DistillBERT for question-answering on the SQuAD dataset. We refer to Appendix C for details.
6.1 Gradient Estimation with Fixed Weights
Synthetic data.
In the first experiment, we verify the hypothesis that the choice of $\mathcal{D}$ matters for problem (4) – in particular, we show that estimating the median is more stable when the noise distribution is heavy-tailed.
We consider the problem of estimating a fixed vector $\bm{\hat{g}}$ from the oracle $\bm{g}\sim\bm{\hat{g}}+\bm{\xi}$ with $\bm{\xi}\sim\mathcal{P}$, under varying degree of heavy-tailedness of $\mathcal{P}$. For this purpose, we choose the standard, unskewed $\alpha$-stable distribution $\mathcal{P}_{\alpha}$.
Setup. We generate $\bm{\hat{g}}\in\mathbb{R}^{d}$ with $d=10$ where each component is generated i.i.d standard Gaussian. In each iteration, a sample $\bm{g}_{t}$ is generated as follows: each coordinate of $(\bm{g}_{t})_{i}$ is (independently) sampled from $\mathcal{P}_{\alpha}$ with location $(\hat{\bm{g}})_{i}$ and varying values for $0<\alpha\leq 2$. Importantly, the median of $\mathcal{P}_{\alpha}$ is equal to the location (i.e. $\bm{\hat{g}}$) for all values of $\alpha$. Recall that for $\alpha\leq 1$ the mean of $\mathcal{P}_{\alpha}$ is not defined, and otherwise equal to the median.
We run the SPP method (9) for $\mathcal{D}\in\{\tfrac{1}{2}\|\cdot\|^{2},\|\cdot\|_{2},\|\cdot\|_{1}\}$, with a fixed step size $\tau=0.01$ and track the relative $\ell_{2}$-error $\frac{\|\bm{m}_{t}-\bm{\hat{g}}\|_{2}}{\|\bm{\hat{g}}\|_{2}}$. The corresponding methods are called momentum, VClip, and CClip. For each method and distribution, we run $1000$ iterations and $50$ different seeds.
Discussion. From Fig. 3 we find that as the noise becomes more heavy-tailed (as $\alpha$ decreases), momentum fails to produce accurate estimates of $\bm{\hat{g}}$. On the other hand, both VClip and CClip are robust to heavy tails, as expected, as we showed that VClip and CClip are estimating the median.
Figure 3: $\tau=0.01$ Left: Final error for varying values of $\alpha$ (from left to right, distributions are more heavy-tailed). Shaded area marks minimal and maximal value over the $50$ independent runs. Right: Convergence plot for all methods for $\tfrac{1}{\alpha}\in[0.5,1.5]$ (higher value of $\tfrac{1}{\alpha}$ corresponds to heavier tails).
Language modelling.
We also run momentum, VClip and CClip in order to estimate the full gradient for the language modeling tasks, where the weights are fixed at initialization. The setup is identical to the one described in (Kunstner et al., 2023), in particular Figure 1 therein. We present the results in Section C.1.
6.2 Training with Online Gradient Estimation
Least squares with heavy-tailed noise.
We consider the noise-perturbed least-squares problem $\min_{\bm{w}\in\mathbb{R}^{d}}\tfrac{1}{2}\|\bm{w}\|^{2}$ and use a gradient oracle given by $\bm{g}\sim\bm{w}+\bm{\xi}\in\mathbb{R}^{d}$, where $\bm{\xi}\sim\mathcal{P}$ is the noise term. We consider two setups:
  1. (S1)
    The $d$ components of $\bm{\xi}$ are i.i.d. distributed according to a standard $\alpha$-stable distribution with location zero and $\alpha=1.1$. Note that $\mathbb{E}_{\bm{\xi}\sim\mathcal{P}}\left[\bm{\xi}\right]=\bm{0}$.
  2. (S2)
    Same as (S1), but the noise is state-dependent, given by $\bm{\xi}_{t}=\sigma\cdot\bm{\xi}$ where $\sigma=\sqrt{1+\|\bm{w}_{t}\|^{2}}$.
  3. (S3)
    $\bm{\xi}$ follows an elliptically contoured $\alpha$-stable distribution (Nolan, 2013) with location zero and $\alpha=1.1$. In particular, the components of $\bm{\xi}$ are not independent.
We run six different methods: on the one hand, the method depicted in Section 3, (5), where in each iteration we draw $n=5$ samples. We run three variations of (5), namely using the $\ell_{1}$- and $\ell_{2}$-sample median as well as the sample mean for $\bm{m}_{t}$. The $\ell_{2}$-median is approximately computed using the method proposed in (Vardi & Zhang, 2000).
(a) Setting (S1)
(b) Setting (S2)
(c) Setting (S3)
Figure 4: Least squares with heavy-tailed noise that is independent over components (left), state-dependent (middle), and where components of noise are dependent (right). All methods in grey use $n=5$ samples per iteration, all others only one. Note that for (S2), the sample mean immediately diverges, and hence does not appear in the plot. Shaded area depicts minimum and maximum value over 50 repetitions.
[…]
Discussion. Fig. 4 shows that both $\ell_{1}$- and $\ell_{2}$-sample median attain the smallest objective, while the sample mean does not converge as predicted by our theory (see discussion after Proposition 3.3). From the online methods, CClip performs best in the scenario where the noise is componentwise independent, while VClip is better in the other case. We find that both VClip and CClip reach much smaller loss values than clipping the stochastic gradient directly, and thus clipping only the increment is better suited here. We also find that the $\ell_{1}$-sample median can deal with state-dependent noise (S2) as predicted by our theory.
Language modeling.
We now consider (17), that is training while simultaneously estimating the gradient, for the three language modelling tasks. We compare SGD-M, VClip and CClip, and clipped-SGD. We also add Adam as a baseline, but focus on the comparison among the SGD-type methods. For all methods, we tune the learning rate on a $\log_{10}$-scaled grid (reported in Table 1), displaying the one that attained smallest final training loss (averaged over $5$ seeds).
[…]
Discussion. Fig. 5 shows that VClip improves over SGD-M for PTB and WikiText-2, and is on par for SQuAD. CClip performs always worse, which might be due to the noise being not componentwise independent. However, VClip can not close the gap to Adam, and also performs worse than clipped-SGD for the WikiText-2 dataset.
(a) PTB
(b) WikiText-2
(c) SQuAD
Figure 5: Training loss for each method, with tuned learning rate. Shaded area depicts minimum and maximum value over 5 seeds.
(a) PTB
(b) WikiText-2
(c) SQuAD
Figure 6: Validation score (measured as perplexity, or F1 score) for each method, with tuned learning rate. Shaded area depicts minimum and maximum value over 5 seeds.
7 Conclusion
We provide an analysis of SGD using the median of sampled gradients, and showed convergence guarantees under potentially heavy-tailed and state-dependent noise. In contrast, under the same assumption, SGD with the sample mean does not converge. We further provide a new approach to compute medians using the stochastic proximal point method. Our final algorithm proposes to use online estimates of the median gradient, requiring only one sample per iteration. We observe that this general framework intersects with many existing works on robust stochastic optimization, where the connection with median estimation was not known previously.
APPENDIX
The Appendix is organized as follows.
  • In Appendix A we provide missing proofs for the sample median method gradient descent (SMGD) presented in Section 3.
  • In Appendix B, we provide additional formulas and derivations for the problem of approximating the median, studied in Section 4.
  • In Appendix C, we provide additional plots and supplementary information on numerical experiments.
Appendix A Missing Proofs for Section 3
A.1 On the Assumptions on the Sample Median
We first state a simple fact about Lipschitz functions:
Lemma A.1.
If Assumption 3.1 holds, and if $\ell$ is Lipschitz continuous, then without loss of generality we can assume that $\delta_{2}=\sigma_{2}=0$.
Proof.
If $\ell$ is $G$-Lipschitz, then for every subgradient $\bm{u}\in\partial\ell(\bm{w})$ we have $\|\bm{u}\|\leq G$, so one can replace $\delta_{1}^{2}$ with $\delta_{1}^{2}+\delta_{2}G^{2}$ and $\sigma_{1}^{2}$ with $\sigma_{1}^{2}+\sigma_{2}G^{2}$. ∎
In the following proof of Proposition 3.3 we will need an equivalent but more compact formulation of Assumption 3.1.
Lemma A.2.
Let $\bm{m}$ be a random variable in $\mathbb{R}^{d}$, and consider the following property:
$\mathbb{E}\left[\|\bm{m}-\nabla\ell(\bm{w})\|^{2}\right]\leq\nu_{1}^{2}+\nu_{2}\|\nabla\ell(\bm{w})\|^{2}.$ (20)
If (6) holds then (20) holds with $\nu_{1}^{2}=\delta_{1}^{2}+\sigma_{1}^{2}$ and $\nu_{2}=\delta_{2}+\sigma_{2}$. If (20) holds, then (6) holds with $\delta_{1}=\sigma_{1}=\nu_{1}$ and $\delta_{2}=\sigma_{2}=\nu_{2}$.
Proof.
Expanding the squares, we get
$\mathbb{E}\left[\|\bm{m}-\nabla\ell(\bm{w})\|^{2}\right]=\mathbb{E}\left[\|\bm{m}-\mathbb{E}\left[\bm{m}\right]\|^{2}\right]+\|\mathbb{E}\left[\bm{m}\right]-\nabla\ell(\bm{w})\|^{2},$
where we used that $\mathbb{E}\left[\langle\bm{m}-\mathbb{E}\left[\bm{m}\right],\mathbb{E}\left[\bm{m}\right]-\nabla\ell(\bm{w})\rangle\right]=0$. Now, the first implication is trivial, and the second follows from the positivity of the involved terms. ∎
We now verify that Assumption 3.1 holds true for the $\ell_{1}$-sample median, provided that the noise on the subgradient is linearly depending on an $\alpha$-stable distribution, and that the linear coefficients grow with the norm of the subgradient.
Lemma A.3.
Let $n\in\mathbb{N}$ be odd and $n\geq 3$. Suppose that the subgradient oracles $\bm{g}^{(i)}=\nabla\ell(\bm{w})+\bm{\Sigma}(\bm{w})\bm{\zeta}^{(i)}$ verify Assumption 3.5, for all $i=1,\dots,n$ and set
$\bm{m}=\texttt{median}_{\ell_{1},i\in[n]}\left(\bm{g}^{(i)}(\bm{w})\right).$
Then, Assumption 3.1 holds with $\delta_{1}=\delta_{2}=0$, and $\sigma_{1}^{2}=C_{n}c_{1}$, $\sigma_{2}=C_{n}c_{2}$, for some constant $C_{n}=\mathcal{O}(1/n)$.
Proof.
Let us denote the noise vector by $\bm{\xi}^{(i)}=\bm{\Sigma}(\bm{w})\bm{\zeta}^{(i)}$. By the stability property of $\alpha$-stable distributions (Nolan, 2020, Proposition 1.3), we have that
${\xi}_{k}^{(i)}:=\bm{\xi}_{k}^{(i)}(\bm{w})=\sum_{j=1}^{d}[\bm{\Sigma}(\bm{w})]_{kj}\bm{\zeta}^{(i)}_{j}=^{\mathrm{d}}\|\bm{\Sigma}(\bm{w})_{k}\|_{\alpha}z,$ (21)
where $=^{\mathrm{d}}$ denotes equality in distribution, $z$ is a standard symmetric $\alpha$-stable random variable in $\mathbb{R}$, and $\bm{\Sigma}(\bm{w})_{k}$ denotes the $k$-th row of $\bm{\Sigma}(\bm{w})$. Observe that from our Assumption 3.5 we can write that
$\sum_{i=1}^{d}\|\bm{\Sigma}(\bm{w})_{i}\|_{\alpha}^{2}\leq\rho_{\alpha}^{2}\sum_{i=1}^{d}\|\bm{\Sigma}(\bm{w})_{i}\|_{2}^{2}=\rho_{\alpha}^{2}\|\bm{\Sigma}(\bm{w})\|_{F}^{2}\leq c_{1}^{\prime}+c_{2}^{\prime}\|\nabla\ell(\bm{w})\|^{2},\quad\mbox{for all }\bm{w}\in\mathbb{R}^{d}.$ (22)
by setting $c_{1}^{\prime}=\rho_{\alpha}^{2}c_{1}$ and $c_{2}=\rho_{\alpha}^{2}c_{2}$, and where $\rho_{\alpha}>0$ is the best constant such that $\|\cdot\|_{\alpha}\leq\rho_{\alpha}\|\cdot\|_{2}$, which is $\rho_{\alpha}=d^{\tfrac{2-\alpha}{2\alpha}}$ is $\alpha\leq 2$, and $\rho_{\alpha}=1$ otherwise.
We start by computing the bias term that is required by Assumption 3.1.
$\|\mathbb{E}\left[\bm{m}\right]-\nabla\ell(\bm{w})\|^{2}=$ $\|\mathbb{E}\left[\texttt{median}_{\ell_{1},i\in[n]}\left(\nabla\ell(\bm{w})+\bm{\xi}^{(i)}(\bm{w})\right)\right]-\nabla\ell(\bm{w})\|^{2}$
$=$ $\|\mathbb{E}\left[\texttt{median}_{\ell_{1},i\in[n]}\left(\bm{\xi}^{(i)}(\bm{w})\right)\right]\|^{2}$
where we used the fact that
$\texttt{median}_{\ell_{1},i\in[n]}\left(\nabla\ell(\bm{w})+\bm{\xi}^{(i)}(\bm{w})\right)=\nabla\ell(\bm{w})+\texttt{median}_{\ell_{1},i\in[n]}\left(\bm{\xi}^{(i)}(\bm{w})\right).$
On the other hand, the $\ell_{1}$-median boils down to computing component-wise medians, hence we have that
$\mathbb{E}\left[\texttt{median}_{\ell_{1},i\in[n]}\left(\bm{\xi}^{(i)}(\bm{w})\right)\right]=\begin{bmatrix}\mathbb{E}\left[\texttt{median}_{i\in[n]}\left(\xi_{1}^{(i)}\right)\right]\\ \vdots\\ \mathbb{E}\left[\texttt{median}_{i\in[n]}\left(\xi_{d}^{(i)}\right)\right]\end{bmatrix},$
where $\texttt{median}_{i\in[n]}(\xi_{k}^{(i)})=\underset{m\in\mathbb{R}}{\operatorname{argmin}}\;\frac{1}{n}\sum_{i=1}^{n}|m-\xi_{k}^{(i)}|$ is the one dimensional sample median estimator.
Now, let us compute $\mathbb{E}\left[\texttt{median}_{i\in[n]}\left(\xi^{(i)}_{k}\right)\right]$ for some $k\in\{1,\dots,d\}$. Denoting $r=(n-1)/2\in\mathbb{N}$ and by using the probability density function of the sample median (cf. (Maritz & Jarrett, 1978)), we have that:
$\mathbb{E}\left[\texttt{median}_{i\in[n]}\left(\xi^{(i)}_{k}\right)\right]=$ $\frac{n!}{r!r!}\underbrace{\int_{\mathbb{R}}x\left(F_{\bm{w},k}(x)(1-F_{\bm{w},k}(x))\right)^{r}p_{\bm{w},k}(x)\mathrm{d}x}_{=:\mathrm{A}},$
where $F_{\bm{w},k}$ and $p_{\bm{w},k}(x)$ are respectively the cumulative distribution function (cdf) and the probability density function (pdf) of $\xi_{k}^{(i)}\sim\mathcal{P}_{\alpha,\|\bm{\Sigma}_{k}(\bm{w})\|_{\alpha}}$ that is a symmetric $\alpha$-stable random variable with scale $\|\bm{\Sigma}_{k}(\bm{w})\|_{\alpha}$. As $p_{\bm{w},k}$ is continuous and symmetric around zero, we have $p_{\bm{w},k}(x)=p_{\bm{w},k}(-x)$ and $(1-F_{\bm{w},k}(x))=F_{\bm{w},k}(-x)$ for all $x\in\mathbb{R}$. This gives us:
$\mathrm{A}=$ $\int_{0}^{\infty}x\left(F_{\bm{w},k}(x)F_{\bm{w},k}(-x)\right)^{r}p_{\bm{w},k}(x)\mathrm{d}x+\int_{-\infty}^{0}x\left(F_{\bm{w},k}(x)F_{\bm{w},k}(-x)\right)^{r}p_{\bm{w},k}(x)\mathrm{d}x$
$=$ $\int_{0}^{\infty}x\left(F_{\bm{w},k}(x)F_{\bm{w},k}(-x)\right)^{r}p_{\bm{w},k}(x)\mathrm{d}x-\int_{0}^{\infty}x\left(F_{\bm{w},k}(x)F_{\bm{w},k}(-x)\right)^{r}p_{\bm{w},k}(x)\mathrm{d}x$
$=$ $0$
Hence, we obtain that
$\mathbb{E}\left[\texttt{median}_{\ell_{1},i\in[n]}\left(\bm{\xi}^{(i)}(\bm{w})\right)\right]=0.$
This shows that the first line of Assumption 3.1 holds with $\delta_{1}=\delta_{2}=0$.
We now proceed to the estimation of the variance. By using similar arguments, we have that:
$\mathbb{E}\left[\|\bm{m}-\mathbb{E}\left[\bm{m}\right]\|^{2}\right]=\;$ $\mathbb{E}\left[\|\texttt{median}_{\ell_{1},i\in[n]}\left(\bm{\xi}^{(i)}(\bm{w})\right)\|^{2}\right]$
=\;= $\sum_{k=1}^{d}\mathbb{E}\left[\left(\texttt{median}_{i\in[n]}\left(\xi_{k}^{(i)}\right)\right)^{2}\right]$
=\;= $\sum_{k=1}^{d}\frac{n!}{r!r!}\mathbb{E}\left[\Bigl{(}\xi_{k}^{(i)}\Bigr{)}^{2}\Biggl{(}F_{\bm{w},k}\Bigl{(}\xi_{k}^{(i)}\Bigr{)}\Bigl{(}1-F_{\bm{w},k}\Bigl{(}\xi_{k}^{(i)}\Bigr{)}\Bigr{)}\Biggr{)}^{r}\right],$ (23)
where (23) follows from (Maritz & Jarrett, 1978, Equation 2.1). By using (21) with (22) and denoting the cdf of the standard symmetric $\alpha$-stable distribution by $F_{z}$, we further compute that:
$\mathbb{E}\left[\|\bm{m}-\mathbb{E}\left[\bm{m}\right]\|^{2}\right]=\;$ $\frac{n!}{r!r!}\mathbb{E}\left[z^{2}\Biggl{(}F_{z}(z)(1-F_{z}(z)\Biggr{)}^{r}\right]\sum_{k=1}^{d}\|\bm{\Sigma}_{k}(\bm{w})\|_{\alpha}^{2}$
[…]
=:absent:=:\;= : $C_{n}\left(c_{1}^{\prime}+c_{2}^{\prime}\|\nabla\ell(\bm{w})\|^{2}\right).$
This shows that the second line of Assumption 3.1 holds with $\sigma_{1}^{2}=C_{n}\rho_{\alpha}^{2}c_{1}$ and $\sigma_{2}=C_{n}\rho_{\alpha}^{2}c_{2}$.
Finally we observe that, by again (Maritz & Jarrett, 1978, Equation 2.1), $C_{n}$ is the variance of the sample median of a set of i.i.d. standard symmetric $\alpha$-stable variables $\{z_{1},\dots,z_{n}\}$, where $z_{i}=^{\mathrm{d}}z$. As $n\to\infty$, it is well-known that the sample median is asymptotically normal with variance $1/(4nf^{2}(0))$, where $f$ denotes the pdf of $z$ (see (Maritz & Jarrett, 1978)). Hence we conclude that $C_{n}=\mathcal{O}(1/n)$ as $n\to\infty$. This concludes the proof. ∎
A.2 Rates for SMGD for Smooth Nonconvex Functions: Proof of Proposition 3.3
The following proposition is an application of biased SGD results to the SMGD algorithm with errors. It is possible to show that Assumptions 3.1 and 3.2 both imply Assumption 9 from (Demidovich et al., 2023). So from a qualitative point of view the result below is the same that what we would obtain by using Theorem 3 from (Demidovich et al., 2023). But doing so would produce a bound with worse constants, so we prefer to provide a proof exploiting directly our assumptions.
See 3.3
Proof.
In this proof we note $\mathbb{E}_{t}$ for the expectation taken conditionally to the filtration generated by $\bm{w}_{0},\dots\bm{w}_{t}$. In particular, we will have from our Assumption 3.2 on the errors together with the tower rule that $\mathbb{E}_{t}\|\bm{e}_{t}\|^{2}=\mathbb{E}_{t}\mathbb{E}\left[\|\bm{e}_{t}\|^{2}\ |\ \bm{g}_{t}^{(1)},\dots,\bm{g}_{t}^{(n)}\right]\leq\varepsilon^{2}$.
Smoothness of $\ell$ implies
$\ell(\bm{w}_{t+1})-\ell(\bm{w}_{t})-\langle\nabla\ell(\bm{w}_{t}),\bm{w}_{t+1}-\bm{w}_{t}\rangle\leq\frac{L}{2}\|\bm{w}_{t+1}-\bm{w}_{t}\|^{2}.$
If we note $\bm{d}_{t}:=\bm{m}_{t}+\bm{e}_{t}$, thus $\bm{w}_{t+1}-\bm{w}_{t}=-\eta\bm{d}_{t}$. This implies
$\ell(\bm{w}_{t+1})-\ell(\bm{w}_{t})+\eta\langle\nabla\ell(\bm{w}_{t}),\bm{d}_{t}\rangle\leq\frac{L\eta^{2}}{2}\|\bm{d}_{t}\|^{2}.$
Let us bound the two terms depending on $\bm{d}_{t}$. First, take expectation conditioned to $\mathcal{F}_{t}$ (which we will denote $\mathbb{E}_{t}$) to write
$\mathbb{E}_{t}\left[\langle\nabla\ell(\bm{w}_{t}),\bm{d}_{t}\rangle\right]$ $=$ $\langle\nabla\ell(\bm{w}_{t}),\mathbb{E}_{t}\left[\bm{d}_{t}\right]\rangle=\frac{1}{2}\|\nabla\ell(\bm{w}_{t})\|^{2}+\frac{1}{2}\|\mathbb{E}_{t}\left[\bm{d}_{t}\right]\|^{2}-\frac{1}{2}\|\mathbb{E}_{t}\left[\bm{d}_{t}\right]-\nabla\ell(\bm{w}_{t})\|^{2}$
$\geq$ $\frac{1}{2}\|\nabla\ell(\bm{w}_{t})\|^{2}-\frac{1}{2}\|\mathbb{E}_{t}\left[\bm{d}_{t}\right]-\nabla\ell(\bm{w}_{t})\|^{2}.$
Using (6) and $\|\mathbb{E}_{t}\left[\bm{e}_{t}\right]\|^{2}\leq\mathbb{E}_{t}\left[\|\bm{e}_{t}\|^{2}\right]\leq\varepsilon^{2}$ due to Jensen’s inequality, we can bound
$\|\mathbb{E}_{t}\left[\bm{d}_{t}\right]-\nabla\ell(\bm{w}_{t})\|^{2}\leq 2\|\mathbb{E}_{t}\left[\bm{m}_{t}\right]-\nabla\ell(\bm{w}_{t})\|^{2}+2\|\mathbb{E}_{t}\left[\bm{e}_{t}\right]\|^{2}\leq 2\delta_{1}^{2}+2\delta_{2}\|\nabla\ell(\bm{w}_{t})\|^{2}+2\varepsilon^{2}.$
Second, we use Young’s inequality and our assumptions to bound $\|\bm{d}_{t}\|^{2}$:
$\mathbb{E}_{t}\left[\|\bm{d}_{t}\|^{2}\right]$$\leq$ $2\mathbb{E}_{t}\left[\|\bm{e}_{t}+\bm{m}_{t}-\nabla\ell(\bm{w}_{t})\|^{2}\right]+2\|\nabla\ell(\bm{w}_{t})\|^{2}$
$\leq$ $4\mathbb{E}_{t}\left[\|\bm{e}_{t}\|^{2}\right]+4\mathbb{E}_{t}\left[\|\bm{m}_{t}-\nabla\ell(\bm{w}_{t})\|^{2}\right]+2\|\nabla\ell(\bm{w}_{t})\|^{2}$
$\leq$ $4\varepsilon^{2}+4(\delta_{1}^{2}+\sigma_{1}^{2})+(2+4\delta_{2}+4\sigma_{2})\|\nabla\ell(\bm{w}_{t})\|^{2},$
where in the last inequality we used that (6) implies (20) (cf. Lemma A.2). Combine all the above inequalities to get
$\mathbb{E}_{t}\left[\ell(\bm{w}_{t+1})\right]-\ell(\bm{w}_{t})+\tfrac{\eta}{2}\|\nabla\ell(\bm{w}_{t})\|^{2}-\tfrac{\eta}{2}(2\delta_{1}^{2}+2\delta_{2}\|\nabla\ell(\bm{w}_{t})\|^{2}+\varepsilon^{2})$
$\hskip 8.61108pt\leq\frac{L\eta^{2}}{2}\big{(}4(\varepsilon^{2}+\delta_{1}^{2}+\sigma_{1}^{2})+(2+4\delta_{2}+4\sigma_{2})\|\nabla\ell(\bm{w}_{t})\|^{2}\big{)}.$
Rearranging yields
$\mathbb{E}_{t}\left[\ell(\bm{w}_{t+1})\right]-\ell(\bm{w}_{t})$
$\hskip 12.91663pt\leq-\eta\big{[}\tfrac{1}{2}-\delta_{2}-L\eta(1+2\sigma_{2}+2\delta_{2})\big{]}\|\nabla\ell(\bm{w}_{t})\|^{2}+\eta(\delta_{1}^{2}+\varepsilon^{2})+2L\eta^{2}(\varepsilon^{2}+\delta_{1}^{2}+\sigma_{1}^{2}).$
If $\delta_{2}\leq\tfrac{1}{8}$ and using $\eta\leq\frac{1}{8L(1+2\sigma_{2}+2\delta_{2})}$, after taking full expectation, we obtain
$\mathbb{E}\left[\ell(\bm{w}_{t+1})\right]-\mathbb{E}\left[\ell(\bm{w}_{t})\right]$$\leq$ $-\tfrac{\eta}{4}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]+\eta(\delta_{1}^{2}+\varepsilon^{2})+2L\eta^{2}(\varepsilon^{2}+\delta_{1}^{2}+\sigma_{1}^{2}).$
Sum over $t=0,\dots,T-1$, multiply by $\frac{4}{\eta T}$ and use $\min\limits_{t=0,\dots,T-1}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]\leq\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]$ to finally obtain
$\min\limits_{t=0,\dots,T-1}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]\leq\frac{4(\ell(\bm{w}_{0})-\inf\ell)}{\eta T}+8L\eta(\varepsilon^{2}+\delta_{1}^{2}+\sigma_{1}^{2})+4(\delta_{1}^{2}+\varepsilon^{2}).$
A.3 Rates for Biased SGD for Convex Lipschitz Functions
In this section we prove another result about the complexity of biased SGD, that is SGD with biased estimators of the subgradient. As far as we know, biased SGD was only studied in the case of strongly convex functions or smooth nonconvex functions (Demidovich et al., 2023), or in the particular case of random projections. In that regard, the next result is a new contribution to the analysis of biased SGD.
Before stating our result, we recall a few standard definitions given that we are working with a nonsmooth convex function. Given a convex function $\ell:\mathbb{R}^{d}\to\mathbb{R}$, we say that $\bm{u}$ is a subgradient of $\ell$ at $\bm{w}$ if
$(\forall\bm{w}^{\prime}\in\mathbb{R}^{d})\ell(\bm{w}^{\prime})-\ell(\bm{w})-\langle\bm{u},\bm{w}^{\prime}-\bm{w}\rangle\geq 0.$
We note $\partial\ell(\bm{w})$ the set of all subgradients of $\ell$ at $\bm{w}$, that is called the subdifferential of $\ell$ at $\bm{w}$.
Theorem A.4.
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be a convex $G$-Lipschitz function such that ${\rm{argmin}}~{}\ell\neq\emptyset$. Let $\eta>0$, and note $D:=\|\bm{w}_{0}-\bm{w}_{*}\|$ for some $\bm{w}_{*}\in{\rm{argmin}}~{}\ell$. Consider the iterates generated by a biased Stochastic Subgradient Descent method
$\bm{w}_{t+1}=\bm{w}_{t}-\eta\bm{g}_{t},$
and assume that the estimator $\bm{g}_{t}$ has a uniformly finite bias and variance (we denote by $\mathcal{F}_{t}$ the filtration induced by $\bm{w}_{0},\dots,\bm{w}_{t}$):
$\|\mathbb{E}\left[\bm{g}_{t}|\mathcal{F}_{t}\right]-\bm{u}_{t}\|\leq\delta,\quad\mathbb{E}\left[\|\bm{g}_{t}-\mathbb{E}\left[\bm{g}_{t}|\mathcal{F}_{t}\right]\|^{2}|\mathcal{F}_{t}\right]\leq\sigma^{2},$
where $\bm{u}_{t}$ is any subgradient of $\ell$ at $\bm{w}_{t}$. Then, for every $T\geq 1$
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{2D^{2}}{\eta T}+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\eta\frac{\delta^{2}T}{4}$
where $\bar{\bm{w}}_{T}$ is an average of the first $T$ iterates defined by $\bar{\bm{w}}_{T}=\frac{1}{\sum_{t=0}^{T-1}\theta^{t}}\sum_{t=0}^{T-1}\theta^{t}\bm{w}_{t}$, $\theta=\frac{T}{T+2}$. In particular, if we take a constant step size equal to
$\eta=\frac{2\sqrt{2}D}{\sqrt{2(\sigma^{2}+(\delta+G)^{2})T+\delta^{2}T^{2}}}\quad\text{(resp. $\eta=\frac{1}{\delta T+\sqrt{T}}$ )}$
then we can guarantee that
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq 2D\sqrt{\frac{\sigma^{2}+(\delta+G)^{2}}{T}+\frac{\delta^{2}}{2}}\quad\text{(resp. $\frac{2D^{2}+\sigma^{2}+\delta^{2}+G^{2}}{\sqrt{T}}+(2D^{2}+1)\delta$ ).}$.
Proof.
Let $w_{*}\in{\rm{argmin}}~{}\ell$ and $T\geq 1$ be fixed. Develop the squares and use the definition of the algorithm to write
$\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}$ $=$ $\|\bm{w}_{t}-\bm{w}_{*}\|^{2}+2\langle\bm{w}_{t+1}-\bm{w}_{t},\bm{w}_{t}-\bm{w}_{*}\rangle+\|\bm{w}_{t+1}-\bm{w}_{t}\|^{2}$
$=$ $\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta\langle\bm{g}_{t},\bm{w}_{t}-\bm{w}_{*}\rangle+\eta^{2}\|\bm{g}_{t}\|^{2}.$
Let $\bar{\bm{g}}_{t}$ be the projection of $\mathbb{E}\left[\bm{g}_{t}|\mathcal{F}_{t}\right]$ onto $\partial\ell(\bm{w}_{t})$, the subdifferential of $\ell$ at $\bm{w}_{t}$. This means that from our assumption on the bias of $\bm{g}_{t}$ we will have
$\|\mathbb{E}\left[\bm{g}_{t}|\mathcal{F}_{t}\right]-\bar{\bm{g}}_{t}\|\leq\|\mathbb{E}\left[\bm{g}_{t}|\mathcal{F}_{t}\right]-\bm{u}_{t}\|\leq\delta.$
Then, after taking conditional expectation (denoted by $\mathbb{E}_{t}\left[\cdot\right]$ instead of $\mathbb{E}\left[\cdot|\mathcal{F}_{t}\right]$) we can write
$\mathbb{E}_{t}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]$ $=$ $\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta\langle\mathbb{E}_{t}\left[\bm{g}_{t}\right],\bm{w}_{t}-\bm{w}_{*}\rangle+\eta^{2}\mathbb{E}_{t}\left[\|\bm{g}_{t}\|^{2}\right]$
$=$ $\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta\langle\bar{\bm{g}}_{t},\bm{w}_{t}-\bm{w}_{*}\rangle+2\eta\langle\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right],\bm{w}_{t}-\bm{w}_{*}\rangle+\eta^{2}\mathbb{E}_{t}\left[\|\bm{g}_{t}\|^{2}\right]$
$\leq$ $\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta(\ell(\bm{w}_{t})-\inf\ell)+2\eta\langle\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right],\bm{w}_{t}-\bm{w}_{*}\rangle+\eta^{2}\mathbb{E}_{t}\left[\|\bm{g}_{t}\|^{2}\right],$
where in the last inequality we used the convexity of $\ell$ together with the fact that $\bar{\bm{g}}_{t}\in\partial\ell(\bm{w}_{t})$. The last term can be bounded by using our assumptions on the estimator $\bm{g}_{t}$, and the fact that $\ell$ has bounded subgradients:
$\mathbb{E}_{t}\left[\|\bm{g}_{t}\|^{2}\right]=\mathbb{E}_{t}\left[\|\bm{g}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right]\|^{2}\right]+\|\mathbb{E}_{t}\left[\bm{g}_{t}\right]\|^{2}\leq\sigma^{2}+(\|\mathbb{E}_{t}\left[\bm{g}_{t}\right]-\bar{\bm{g}}_{t}\|+\|\bar{\bm{g}}_{t}\|)^{2}\leq\sigma^{2}+(\delta+G)^{2}.$
Injecting this in the previous inequality we obtain
$\mathbb{E}_{t}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]\leq\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta(\ell(\bm{w}_{t})-\inf\ell)+2\eta\langle\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right],\bm{w}_{t}-\bm{w}_{*}\rangle+\eta^{2}\sigma^{2}+\eta^{2}(\delta+G)^{2}.$
Now we introduce a parameter $\eta>0$, and we use Young’s inequality together with our assumption on the bias of $\bm{g}_{t}$ to write
$2\langle\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right],\bm{w}_{t}-\bm{w}_{*}\rangle\leq\frac{1}{\eta}\|\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right]\|^{2}+\eta\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\leq\eta^{-1}\delta^{2}+\eta\|\bm{w}_{t}-\bm{w}_{*}\|^{2}.$
Inject this bound in our main inequality
$\mathbb{E}_{t}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]\leq(1+\eta\eta)\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta(\ell(\bm{w}_{t})-\inf\ell)+\eta^{2}\sigma^{2}+\eta^{2}(\delta+G)^{2}+\eta^{-1}\delta^{2},$
and after reorganizing the terms, dividing by $2\eta$ and taking expectation we finally obtain
$\mathbb{E}\left[\ell(\bm{w}_{t})-\inf\ell\right]\leq\frac{1+\eta\eta}{2\eta}\mathbb{E}\left[\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\right]-\frac{1}{2\eta}\mathbb{E}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\eta^{-1}\frac{\delta^{2}}{2}.$
We are now ready to use a weighted telescopic sum argument. First, define $\theta:=\frac{1}{1+\eta\eta}$ and $\rho_{t}:=\theta^{t}$. Second, multiply our inequality by $\rho_{t+1}$ and then sum for $t=0,\dots,T-1$. Observe that the term $(1+\eta\eta)\rho_{t}$ is equal to $\rho_{t}$ due to our definition, which means that we have a telescoping sum where the terms $\rho_{t}\mathbb{E}\left[\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\right]$ will cancel each other. Third, divide by $\sum_{t=0}^{T-1}\rho_{t+1}$ so to obtain
$\frac{1}{\sum_{t=0}^{T-1}\rho_{t+1}}\sum_{t=0}^{T-1}\rho_{t+1}\mathbb{E}\left[\ell(\bm{w}_{t})-\inf\ell\right]\leq\frac{\rho_{0}}{2\eta\sum_{t=0}^{T-1}\rho_{t+1}}\mathbb{E}\left[\|\bm{w}_{0}-\bm{w}_{*}\|^{2}\right]+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\eta^{-1}\frac{\delta^{2}}{2}.$
Now define $\bar{\bm{w}}_{T}:=\frac{1}{\sum_{t=0}^{T-1}\rho_{t+1}}\sum_{t=0}^{T-1}\rho_{t+1}\bm{w}_{t}=\frac{1}{\sum_{t=0}^{T-1}\theta^{t}}\sum_{t=0}^{T-1}\theta^{t}\bm{w}_{t}$, and use Jensen’s inequality to get
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{\|\bm{w}_{0}-\bm{w}_{*}\|^{2}}{2\eta\sum_{t=0}^{T-1}\rho_{t+1}}+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\eta^{-1}\frac{\delta^{2}}{2}.$
We will now simplify this upper bound, so we can later make an appropriate choice of $\alpha$ and $\eta$ leading to the desired complexity bound. We will focus now on the geometric sum, and start by using the definition of $\rho_{t}$ to write
$\eta\sum_{t=0}^{T-1}\rho_{t+1}=\eta\theta\sum_{t=0}^{T-1}\theta^{t}=\eta\frac{\theta}{1-\theta}(1-\theta^{T}).$
On the one hand, we see that from the definition of $\theta$ we have
$\eta\frac{\theta}{1-\theta}=\frac{\eta}{1+\eta\eta}\frac{1}{1-\frac{1}{1+\eta\eta}}=\frac{\eta}{\eta\eta}=\eta^{-1}.$
On the other hand, we can guarantee that $(1-\theta^{T})\geq 1/2$ provided that $\eta=\frac{2}{\eta T}$. To see this, first observe that $(1-\theta^{T})\geq 1/2$ is equivalent to $T\geq\frac{\ln(2)}{\ln(\theta^{-1})}$. Since we impose that $\eta\eta=\frac{2}{T}$ then necessarily $\theta=\frac{T}{T+2}\Leftrightarrow\theta^{-1}=1+\frac{2}{T}$. So what we need is $T\geq\frac{\ln(2)}{\ln(1+\frac{2}{T})}$. Now, observe that $\ln(2)\leq 1$. Moreover, it is an exercise (see Lemma A.5 which is deferred to the end of this proof) to verify that for all $x\geq 0$, $\frac{1}{\ln(x+1)}\leq\frac{1}{2}+\frac{1}{x}$. So for our condition to hold it is enough to ask that $T\geq\frac{1}{2}+\frac{T}{2}$, which is equivalent to $T\geq 1$, which is true.
Combining all those bounds together with our new definition for $\eta$, we now have
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{2\|\bm{w}_{0}-\bm{w}_{*}\|^{2}}{\eta T}+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\frac{\eta\delta^{2}T}{4}.$
To simplify the last stage of our analysis, let us note $a=2\|\bm{w}_{0}-\bm{w}_{*}\|^{2}$, $b=\frac{\sigma^{2}+(\delta+G)^{2}}{2}$ and $c=\frac{\delta^{2}}{4}$, so that our bound writes as
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{a}{\eta T}+\eta b+c\eta T.$ (24)
Minimizing the right-hand side is equivalent to solve $\frac{a}{\eta T}=\eta b+c\eta T$, where
$\frac{a}{\eta T}=\eta b+c\eta T\Leftrightarrow\eta^{2}=\frac{a}{T(b+cT)}\Leftrightarrow\eta=\sqrt{\frac{a}{bT+cT^{2}}}\Leftrightarrow\eta=2\sqrt{\frac{2\|\bm{w}_{0}-\bm{w}_{*}\|^{2}}{2(\sigma^{2}+(\delta+G)^{2})T+\delta^{2}T^{2}}}$
With such choice of stepsize, our bound becomes
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{2a}{\eta T}=\frac{2a}{T}\sqrt{\frac{bT+cT^{2}}{a}}=2\sqrt{a}\sqrt{\frac{b}{T}+c}=2\|\bm{w}_{0}-\bm{w}_{*}\|\sqrt{\frac{\sigma^{2}+(\delta+G)^{2}}{T}+\frac{\delta^{2}}{2}}$
and this gives us our first bound. For the second bound, let us take simply $\eta=\frac{1}{\delta T+\sqrt{T}}$ and inject it into (24) to obtain
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]$$\leq$ $\frac{a\delta T+a\sqrt{T}}{T}+\frac{b}{\delta T+\sqrt{T}}+\frac{cT}{\delta T+\sqrt{T}}$
$\leq$ $a\delta+\frac{a}{\sqrt{T}}+\frac{b}{\sqrt{T}}+\frac{c}{\delta}=\frac{a+b}{\sqrt{T}}+a\delta+\frac{c}{\delta}$
$\leq$ $\frac{2D^{2}+\sigma^{2}+\delta^{2}+G^{2}}{\sqrt{T}}+(2D^{2}+1)\delta,$
where in the last inequality we simplified some numerical constants. ∎
Lemma A.5.
For every $x\geq 0$, $\frac{1}{\ln(1+x)}\leq\frac{1}{2}+\frac{1}{x}$.
Proof.
This inequality is equivalent to $\ln(1+x)\geq\frac{2x}{x+2}$, or again $(x+2)\ln(1+x)\geq 2x$. Define $\phi:(-1,+\infty)\to\mathbb{R}$ as $\phi(x)=(x+2)\ln(1+x)$ and compute its derivatives:
$\phi^{\prime}(x)=\ln(1+x)+1+\frac{1}{1+x}\quad\text{ and }\quad\phi^{\prime\prime}(x)=\frac{x}{(1+x)^{2}}.$
We see that $\phi^{\prime\prime}(x)\geq 0$ for all $x\geq 0$, so $\phi$ is convex on $[0,+\infty)$. So we can use the tangent inequality:
$\phi(x)\geq\phi(0)+\phi^{\prime}(0)(x-0)=0+2x=2x.$ (25)
Therefore $\phi(x)\geq 2x$ for all $x\geq 0$, which is what we wanted to prove. ∎
A.4 Rates for SMGD for Convex Lipschitz Functions: Proof of Proposition 3.4
See 3.4
Proof.
In this proof, we note $\mathcal{F}_{t}$ the filtration generated by $\bm{w}_{0},\dots,\bm{w}_{t}$, and we will note $\mathbb{E}_{t}$ to refer to the expectation conditioned on $\mathcal{F}_{t}$.
We want to apply Theorem A.4. Let us denote $\bm{g}_{t}=\bm{m}_{t}+\bm{e}_{t}$, so that the iterates of SMGD verify $\bm{w}_{t+1}=\bm{w}_{t}-\eta\bm{g}_{t}$. We are now going to show that $\bm{g}_{t}$ has uniformly bounded bias and variance. For this we will make use of Assumptions 3.1 and 3.2. Remember that we assumed $\delta_{2}=\sigma_{2}=0$ without loss of generality (we can do this according to Lemma A.1), so we note $\delta$ and $\sigma$ instead of $\delta_{1}$ and $\sigma_{1}$. Also, Assumption 3.2 on the errors together with the tower rule imply that $\mathbb{E}_{t}\left[\|\bm{e}_{t}\|^{2}\right]=\mathbb{E}_{t}\left[\mathbb{E}\left[\|\bm{e}_{t}\|^{2}\ |\ \bm{g}_{t}^{(1)},\dots,\bm{g}_{t}^{(n)}\right]\right]\leq\varepsilon^{2}$. Regarding the bias, we can write
$\|\mathbb{E}_{t}\left[\bm{g}_{t}-\nabla\ell(\bm{w}_{t})\right]\|\leq\|\mathbb{E}_{t}\left[\bm{m}_{t}-\nabla\ell(\bm{w}_{t})\right]\|+\|\mathbb{E}_{t}\left[\bm{e}_{t}\right]\|\leq\delta+\varepsilon,$
where in the last inequality we used Jensen’s inequality to write $\|\mathbb{E}_{t}\left[\bm{e}_{t}\right]\|\leq\sqrt{\mathbb{E}_{t}\left[\|\bm{e}_{t}\|^{2}\right]}$. As for the variance, we write
$\mathbb{E}_{t}\left[\|\bm{g}_{t}-\mathbb{E}_{t}\bm{g}_{t}\|^{2}\right]\leq 2\mathbb{E}_{t}\left[\|\bm{m}_{t}-\mathbb{E}_{t}\left[\bm{m}_{t}\right]\|^{2}\right]+2\mathbb{E}_{t}\left[\|\bm{e}_{t}-\mathbb{E}_{t}\left[\bm{e}_{t}\right]\|^{2}\right]\leq 2\sigma^{2}+8\mathbb{E}_{t}\left[\|\bm{e}_{t}\|^{2}\right]\leq 2\sigma^{2}+8\varepsilon^{2}.$
Using Theorem A.4, we obtain
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{2\|\bm{w}_{0}-\bm{w}_{*}\|^{2}}{\eta T}+\eta\frac{2\sigma^{2}+8\varepsilon^{2}+(\delta+\varepsilon+G)^{2}}{2}+\frac{\eta(\delta+\varepsilon)^{2}T}{4}.$
Setting the step size $\eta=\frac{1}{(\delta+\varepsilon)T+\sqrt{T}}$, we conclude that
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{2D^{2}+2\sigma^{2}+8\varepsilon^{2}+(\delta+\varepsilon)^{2}+G^{2}}{\sqrt{T}}+(2D^{2}+1)(\delta+\varepsilon).$
A.5 Rates for SMGD with Stable Noise: Proof of Corollary 3.6
See 3.6
Proof.
First of all, remember that the $\ell_{1}$-median can be computed exactly, by computing component-wise a one-dimensional median. This allows us to have $\varepsilon=0$ in Assumption 3.2.
Proof of part (i). The result is a direct consequence of Lemma A.3 and Proposition 3.3.
Proof of part (ii). The result is a direct consequence of Lemma A.3 and Proposition 3.4.
This completes the proof. ∎
Appendix B Approximate Computation of Median-like Estimators
B.1 Median-like Estimators : General Considerations
Given a function $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ we consider the problem (4), which writes as
$\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right].$
In what follows, we will mostly consider functions $\mathcal{D}$ satisfying the following assumption.
Assumption B.1.
The function $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, ${\rm{argmin}}~{}\mathcal{D}=\{0\}$ and the problem (4) admits a solution.
Note that assuming $\mathcal{D}$ to be convex and finite implies that it is continuous, therefore measurable, which means that $\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right]$ is well defined. The assumption that ${\rm{argmin}}~{}\mathcal{D}=\{0\}$ guarantees that in the trivial case when the distribution is a Dirac $\delta_{\bm{m}_{0}}$, then the solution to (4) is exactly $\bm{m}_{0}$. The assumption that a solution to (4) exists is easily verified in practice, under the assumption that $\mathcal{D}$ is coercive:
Lemma B.2.
Let $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ be convex. Assume that $\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]<+\infty$. If $\mathcal{D}$ is coercive, i.e. $\lim\limits_{\|m\|\to\infty}\mathcal{D}(\bm{m})=+\infty$, then (4) admits a solution.
Proof.
Let $\phi(\bm{m}):=\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right]$. Since we assume that $\phi(\mathbb{E}\left[\bm{z}\right])=\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]<+\infty$ then $\phi$ is proper. Since $\mathcal{D}$ is convex then $\phi$ is convex, as an expectation of convex functions. Moreover $\mathcal{D}$ is convex and finite, which means that $\mathcal{D}$ is continuous on $\mathbb{R}^{d}$, see e.g. Corollary 8.39 in (Bauschke & Combettes, 2017). Therefore $\phi$ is an expectation of lower semicontinuous functions, which implies that $\phi$ is lower semicontinuous.
Since $\mathcal{D}$ is coercive, convex and continuous, we know that there exists $a\in(0,+\infty)$ and $b\in\mathbb{R}$ (see Proposition 14.16 in (Bauschke & Combettes, 2017)) such that $\mathcal{D}(\cdot)\geq a\|\cdot\|+b$. Therefore
$\phi(\bm{m})$ $=$ $\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right]\geq a\mathbb{E}\left[\|\bm{m}-\bm{z}\|\right]+b$
$\geq$ $a\|\bm{m}-\mathbb{E}\left[\bm{z}\right]\|-a\mathbb{E}\left[\|\bm{z}-\mathbb{E}\left[\bm{z}\right]\|\right]+b$
$\geq$ $a\|\bm{m}-\mathbb{E}\left[\bm{z}\right]\|-a\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]\underset{\|\bm{m}\|\to\infty}{\longrightarrow}+\infty,$
[…]
B.2 Approximations based on Subgradient Updates
Instead of SPP, we could solve (8) by iteratively taking steps of stochastic subgradient descent. If $\mathcal{D}$ is convex, then in each iteration we sample a gradient $\bm{g}_{t}$, and update our current estimate $\bm{m}_{t}$ by
$\bm{m}_{t+1}=\bm{m}_{t}-\tau\bm{u}_{t},\quad\bm{u}_{t}\in\partial\mathcal{D}(\bm{m}_{t}-\bm{g}_{t}).$ (26)
Here, $\tau>0$ is the learning rate, and $\bm{u}_{t}$ is a (convex) subgradient. Since $\mathcal{D}$ could be a non-differentiable function such as the $\ell_{2}$- or $\ell_{1}$-norm, we use subgradients instead of gradients.
Squared $\ell_{2}$-norm.
As a first simple example of (26), let $\mathcal{D}=\tfrac{1}{2}\|\cdot\|_{2}^{2}.$ In this case (26) becomes
$\bm{m}_{t+1}=\bm{m}_{t}-\tau(\bm{m}_{t}-\bm{g}_{t})=(1-\tau)\bm{m}_{t}+\tau\bm{g}_{t}.$ (27)
[…]
$\ell_{1}$-norm.
If we choose $\mathcal{D}=\|\cdot\|_{1}$, update (26) gives
$\bm{m}_{t+1}=\bm{m}_{t}+\tau\operatorname{sgn}(\bm{g}_{t}-\bm{m}_{t}),$
where the $\operatorname{sgn}$-operator can take any value in $[-1,1]$ for coordinates $i\in[d]$ such that $(\bm{g}_{t}-\bm{m}_{t})_{i}=0$. This recovers a version of sign-SGD (Riedmiller & Braun, 1993; Bernstein et al., 2018), where the $\mathrm{sign}$-operation is applied to the increment $\bm{m}_{t}-\bm{g}_{t}$. For example, if we reset $\bm{m}_{t}=0$ in every iteration, we obtain exactly sign-SGD.
$\ell_{2}$-norm.
If we choose $\mathcal{D}=\|\cdot\|_{2}$, and if $\bm{g}_{t}\neq\bm{m}_{t}$, update (26) gives
$\bm{m}_{t+1}=\bm{m}_{t}-\tau\frac{\bm{m}_{t}-\bm{g}_{t}}{\|\bm{m}_{t}-\bm{g}_{t}\|}=\big{(}1-\frac{\tau}{\|\bm{m}_{t}-\bm{g}_{t}\|}\big{)}\bm{m}_{t}+\frac{\tau}{\|\bm{m}_{t}-\bm{g}_{t}\|}\bm{g}_{t}.$
Otherwise, if $\bm{g}_{t}=\bm{m}_{t}$, we can choose any $\bm{u}_{t}$ with $\|\bm{u}_{t}\|\leq 1$ and set $\bm{m}_{t+1}=\bm{m}_{t}-\tau\bm{u}_{t}$; in particular we can set $\bm{m}_{t+1}=\bm{m}_{t}$.
B.3 Approximations based on the Stochastic Proximal Point: General $\mathcal{D}$
Given a proper, closed convex function $\phi:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ and $\tau>0$, we recall the definition of the proximal operator
$\mathrm{prox}_{\tau\phi}(\bm{m}_{0})=\underset{\bm{m}\in\mathbb{R}^{d}}{\rm{argmin}}~{}\phi(\bm{m})+\frac{1}{2\tau}\|\bm{m}-\bm{m}_{0}\|^{2}.$
Applying the SPP algorithm to (8) gives the following iterations
$\begin{cases}\text{Sample $\bm{g}_{t}$ i.i.d. from $\mathcal{G}$}\\ \bm{m}_{t+1}=\mathrm{prox}_{\tau\mathcal{D}(\cdot-\bm{g}_{t})}(\bm{m}_{t}).\end{cases}$ (28)
As we see next, those iterations can be reformulated to simply involve the proximal operator of $\mathcal{D}$:
Proposition B.3.
Let $\mathcal{D}$ verify Assumption B.1. The SPP update (28) for solving (8) can be equivalently written as
$\bm{m}_{t+1}$ $=\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t})$ (29)
$=\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\bm{m}_{t}-\cdot}{\tau})}(\bm{g}_{t}),$ (30)
where $\mathcal{D}^{*}(\bm{m})=\sup_{\bm{y}}\left<\bm{y},\bm{m}\right>-\mathcal{D}(\bm{y})$ is the Fenchel conjugate of $\mathcal{D}$, and $\bm{g}_{t}$ is sampled i.i.d. from $\mathcal{G}$ at each iteration.
Proof.
Here we make use of proximal calculus rules, which can found in most textbooks on convex analysis, such as chapter 6 in (Beck, 2017). Starting from (9), with $\bm{y}=\bar{\bm{y}}+\bm{g}_{t}$, we have
$\begin{split}&\bm{m}_{t+1}=\underset{\bm{y}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathcal{D}(\bm{y}-\bm{g}_{t})+\tfrac{1}{2\tau}\|\bm{y}-\bm{m}_{t}\|^{2}\\ &\hskip 4.30554pt=\bm{g}_{t}+\underset{\bar{\bm{y}}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathcal{D}(\bar{\bm{y}})+\tfrac{1}{2\tau}\|\bar{\bm{y}}-(\bm{m}_{t}-\bm{g}_{t})\|^{2}\\ &\hskip 4.30554pt=\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t}).\end{split}$
This proves (29). As for (30), we combine $(\tau\mathcal{D})^{*}=\tau\mathcal{D}^{*}(\cdot/\tau)$ (Beck, 2017, Thm. 4.14) and the Moreau decomposition theorem (Beck, 2017, Thm. 6.44), to obtain
$\mathrm{prox}_{\tau\mathcal{D}}(\bm{m})=\bm{m}-\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\cdot}{\tau})}(\bm{m})$
in (29) gives
$\bm{m}_{t+1}$ $=\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t})$
$=\bm{g}_{t}+\bm{m}_{t}-\bm{g}_{t}-\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\cdot}{\tau})}(\bm{m}_{t}-\bm{g}_{t})$
$=\bm{m}_{t}-\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\cdot}{\tau})}(\bm{m}_{t}-\bm{g}_{t})$ (31)
Finally, using that for any function $f$ it holds $\underset{\bm{y}}{\operatorname{argmin}}\;f(\bm{y})=\bm{m}+\underset{\bar{\bm{y}}}{\operatorname{argmin}}\;f(\bar{\bm{y}}+\bm{m})$ in (31) gives
$\bm{m}_{t+1}$ $=\bm{m}_{t}-\left(\bm{m}_{t}+\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\cdot+\bm{m}_{t}}{\tau})}(-\bm{g}_{t})\right)$
$=-\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\cdot+\bm{m}_{t}}{\tau})}(-\bm{g}_{t})$
$=\mathrm{prox}_{\tau\mathcal{D}^{*}(\frac{\bm{m}_{t}\cdot-}{\tau})}(\bm{g}_{t})$
where in the final equality we used that if $f(\bm{x})=g(-\bm{x})$ we have that
$\mathrm{prox}_{g}(\bm{m})=-\mathrm{prox}_{f}(-\bm{m}).$
B.4 Approximations based on the Stochastic Proximal Point: Particular Cases of $\mathcal{D}$
$\ell_{p}$-norm.
We now specialize the previous results for when $\mathcal{D}$ is an $\ell_{p}$-norm. See 4.1
Proof.
The proximal operator of the $\ell_{p}$-norm is given by
$\mathrm{prox}_{\tau\|\cdot\|_{p}}(\bm{m})=\bm{m}-\tau\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0})}(\bm{m}/\tau),$
see Example 6.47 in (Beck, 2017). Using this in (29) gives
$\bm{m}_{t+1}$ $=\bm{g}_{t}+\mathrm{prox}_{\tau\|\cdot\|_{p}}(\bm{m}_{t}-\bm{g}_{t})$
$=\bm{m}_{t}-\tau\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0},1)}\big{(}\frac{\bm{m}_{t}-\bm{g}_{t}}{\tau}\big{)}$
$=\bm{m}_{t}+\tau\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0},1)}\big{(}\frac{\bm{g}_{t}-\bm{m}_{t}}{\tau}\big{)}$
$=\bm{m}_{t}+\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0},\tau)}\big{(}\bm{g}_{t}-\bm{m}_{t}\big{)}$
$=\bm{m}_{t}+\left(\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m}_{t},\tau)}\big{(}\bm{g}_{t}\big{)}-\bm{m}_{t}\right)$
$=\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m}_{t},\tau)}\big{(}\bm{g}_{t}\big{)}$
where the third equality we used that
$\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0},1)(\bm{m})}=-\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0},1)(-\bm{m})},$
followed by
$\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m}/\tau,1)}(\bm{y}/\tau)=\frac{1}{\tau}\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m},\tau)}(\bm{y})$
in the fourth equality and
$\mathrm{Proj}_{\mathbb{B}_{q}(\bm{0},\tau)}(\bm{y})=\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m},\tau)}(\bm{y}-\bm{m})-\bm{m}$
in the fifth equality. ∎
$\ell_{1}$-norm.
See 4.3
Proof.
Follows from (10) and that
$\mathrm{Proj}_{\mathbb{B}_{2}(\bm{0},\tau)}\big{(}\bm{v}\big{)}=\frac{\tau\;\bm{v}}{\max\{\tau,\|\bm{v}\|_{2}\}}.$
$\ell_{2}$-norm.
See 4.2
Proof.
Follows directly from (10), by using that
$\;\min\{\max\{\bm{v},-\tau\},\tau\}=\mathrm{Proj}_{\mathbb{B}_{\infty}(\bm{0},\tau)}\big{(}\bm{v}\big{)}.$
Huber function.
The Huber function is given by
$H_{\mu}:\mathbb{R}^{d}\to\mathbb{R},\quad H_{\mu}(\bm{z})=\begin{cases}\frac{1}{2\mu}\|\bm{z}\|^{2}&\quad\|\bm{z}\|\leq\mu,\\ \quad\|\bm{z}\|-\frac{\mu}{2}&\quad\text{else}.\end{cases}$
It operates like the squared loss for arguments with small norm, and like the $\ell_{2}$-norm for arguments with large norms (potentially outliers).
Lemma B.4.
Update (9) with $\mathcal{D}=H_{\mu}$ is given by
$\bm{m}_{t+1}$ $=\bm{g}_{t}+\big{(}1-\tfrac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\big{)}(\bm{m}_{t}-\bm{g}_{t})=\big{(}1-\tfrac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\big{)}\bm{m}_{t}+\tfrac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\bm{g}_{t}.$ (32)
Proof.
Follows from (9) and the fact that the proximal operator of the Huber function (see (Beck, 2017, Example 6.66)) is given by $\mathrm{prox}_{\tau H_{\mu}}(\bm{z})=\Big{(}1-\frac{\tau}{\max\{\|\bm{z}\|,\tau+\mu\}}\Big{)}\bm{z}.$
The estimator (32) has the advantage that it is a relaxation of vectorwise clipping (12), as $01-\frac{\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,\tau+\mu\}}\geq 1-\frac{\tau}{\tau+\mu}>$ 0. However, it comes with the disadvantage that it has two hyperparameters $\tau,~{}\mu$ instead of one.
B.5 Proof of Proposition 4.4: Convergence when $\mathcal{D}$ is a Norm
See 4.4
Proof.
Since we assume that $\mathcal{D}$ is a norm over $\mathbb{R}^{d}$ then it convex and continuous. Moreover, since all norms are equivalent in $\mathbb{R}^{d}$, we have that $\mathcal{D}(\bm{w})\leq C\|\bm{w}\|_{2}$ for some constant $G>0$. Thus $\mathcal{D}$ is a $G$-Lipschitz function. If now we write $\varphi_{\bm{g}}:=\mathcal{D}(\cdot-\bm{g})$, it is clear that it is convex and Lipschitz. Therefore, $\varphi=\mathbb{E}_{\bm{g}}\left[\varphi_{\bm{g}}\right]$ is itself a convex $G$-Lipschitz function. We can now conclude with Theorem 4.4 from (Davis & Drusvyatskiy, 2019). ∎
Here is a corollary stating rates on the expected distance between the approximate sample median and the true sample median, when using the $\ell_{1}$-norm:
Corollary B.5.
In the context of Proposition 4.4, assume that $\mathcal{D}=\|\cdot\|_{1}$ and that $\mathcal{G}$ is a sum of Dirac measures $\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\bm{g}^{(i)}}$. Then we have further
$\mathbb{E}\left[\|\bar{\bm{m}}_{T}-\bm{m}_{*}\|\right]=O\left(\frac{\|\bm{m}_{0}-\bm{m}_{*}\|^{2}}{\tau T}+\tau\right).$
Proof.
In this proof, we use the notation $(\bm{x}_{j})_{j=1,\ldots,d}$ to specify the indivdual components of a vector $\bm{x}\in\mathbb{R}^{d}$. For $\bm{x}\in\mathbb{R}^{d}$, we define the sign operator as follows: let $\operatorname{sgn}(\bm{x})=(\operatorname{sgn}(\bm{x}_{j}))_{j=1,\ldots,d}$, where $\bm{x}_{j}$ is the $j$-th coordinate of $\bm{x}\in\mathbb{R}^{d}$, and where we define for a scalar $t\in\mathbb{R}$ the set-valued operator $\operatorname{sgn}(t)=\{+1\}$ if $t>0$, $\operatorname{sgn}(t)=\{-1\}$ if $t<0$, and $\operatorname{sgn}(t)=[-1,1]$ if $t=0$. Further, given a set $A\subset\mathbb{R}^{d}$ we define $\|A\|_{-}:=\inf\limits_{a\in A}\|a\|$.
To obtain the desired bound, and given that we already have the conclusion of Proposition 4.4, it is enough for us to show that there exists $\mu>0$ such that
$(\forall\bm{m}\in\mathbb{R}^{d})\quad\mu\operatorname{dist}(\bm{m},{\rm{argmin}}~{}\varphi)\leq\varphi(\bm{m})-\inf\varphi.$
It is a standard result from variational analysis that this property [^5] It is sometimes referred to as (weak) sharp minima (Ferris, 1991; Burke & Ferris, 1993), (superlinear)(linear) conditioning (Lemaire, 1992; Cornejo et al., 1997; Lemaire, 1998), or error bound (Lewis & Pang, 1998). is true if and only if the following inequality is satisfied (see Theorem 5.1 in (Cornejo et al., 1997) or Proposition 3.1 in (Lemaire, 1998))
$(\forall\bm{m}\in\operatorname{dom}\partial\varphi)\quad\|\partial\varphi(\bm{m})\|_{-}\geq\mu.$
Our goal now is to prove the above inequality, from which the conclusion will follow. Let us consider $\bm{m}\notin{\rm{argmin}}~{}\varphi$, and use standard calculus to write
$\partial\varphi(\bm{m})=\frac{1}{n}\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})=\frac{1}{n}\sum_{i=1}^{n}\big{(}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})\big{)}_{j=1,\ldots,d}=\frac{1}{n}\Big{(}\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})\Big{)}_{j=1,\ldots,d}.$
Our goal is to show that $\|\partial\varphi(\bm{m})\|_{-}\geq\frac{1}{n}$, which is equivalent to show that $\|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})\|_{-}\geq 1$. Because $\bm{m}\notin{\rm{argmin}}~{}\varphi$, we know that $0\notin\partial\varphi(\bm{m})$ so there must exist a coordinate $j$ such that the $j$-th coordinate of $\partial\varphi(\bm{m})$ does not contain zero, in other words such that $0\notin\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})$. Using the fact that $\|\cdot\|_{2}\geq\|\cdot\|_{\infty}$, we can therefore lower bound $\|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}-{\bm{g}}_{i})\|_{-}\geq|\sum_{i=1}^{n}\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})|_{-}$. Let us denote $s_{i}:=\operatorname{sgn}(\bm{m}_{j}-({\bm{g}}_{i})_{j})$, so that we want to prove that $|\sum_{i=1}^{n}s_{i}|_{-}\geq 1$. Let us now consider a few cases.
  • If $s_{i}=\{+1\}$ or $\{-1\}$ for every $i$, then $\sum_{i=1}^{n}s_{i}$ is a singleton. Furthermore it is a sum of relative numbers, so $\sum_{i=1}^{n}s_{i}\in\mathbb{Z}$. But we also know that $0\notin\sum_{i=1}^{n}s_{i}$, so we conclude that $\sum_{i=1}^{n}s_{i}\in\mathbb{Z}^{*}$, and so that $|\sum_{i=1}^{n}s_{i}|_{-}\geq 1$.
  • If $s_{i}=[-1,+1]$ for every $i$, then we immediately see that $0\in\sum_{i=1}^{n}s_{i}$ which is a contradiction.
  • Otherwise, the $s_{i}$ are combinations of singletons and intervals. Let us note $I=\{i:s_{i}=[-1,+1]\}$ and $I^{\prime}=\{i:s_{i}=\{\pm 1\}\}$, which are not empty by assumption. Let us also note $k\geq 1$ the cardinality of $I$, so that $\sum_{i\in I}s_{i}=[-k,+k]$. We also introduce $s:=\sum_{i\in I^{\prime}}s_{i}\in\mathbb{Z}$. Now we can write
    $|\sum_{i=1}^{n}s_{i}|_{-}=\inf\limits_{t\in\sum_{i=1}^{n}s_{i}}|t|=\inf\limits_{t\in[-k,+k]}|t+s|=d(s;[-k,+k]),$
    where the latter is the distance from the number $s$ to the interval $[-k,+k]$. Now remember that $0\notin\sum_{i=1}^{n}s_{i}$, so this distance cannot be zero, which means that $|s|>k$. Moreover both $s$ and $k$ are rational numbers, so this distance must be greater or equal to $1$. This concludes the proof.
Appendix C Supplementary Information on Experiments
This section provides additional information for the experimental setup.
For easier interpretation, we use the first order approximation of momentum (14), namely $\bm{m}_{t+1}=\left(1-\tau\right)\bm{m}_{t}+\tau\bm{g}_{t}$. That is, we replace $\frac{\tau}{1+\tau}$ with its first-order Taylor approximation around zero, which is equal to $\tau$. Note that $\tau$ is typically small, for example for $\tau=0.01$ we have $\frac{\tau}{1+\tau}\approx 0.0099$. This applies only to the experiments with estimating a fixed vector (or gradient) in Section 6.1.
For the language modeling experiments, all details are identical to (Kunstner et al., 2023), Section A.1. They also provide implementations for all tasks at https://github.com/fKunstner/noise-sgd-adam-sign, which we use. For the training runs for language modeling, we use batch size $256$ for PTB, $320$ for WikiText-2, and $32$ for SQuAD.
C.1 Additional Plots
Here, we show the results for estimating the full gradient of a transformer architecture, with weights fixed at initialization. Fig. 7 shows the results for PTB where we tried three different batch sizes $\{64,256,1024\}$. Fig. 8 shows the results for WikiText-2.
Discussion. From Fig. 7, we observe two phenomena: in the long run, momentum attains the lowest error for estimating the full batch gradient. However, the initial decrease of the error is much faster for CClip, followed by VClip. This is important when using these estimates within a training setup such as (17), where we only do one iteration of gradient estimation, followed by an update of the weight (and hence a change in the full-batch gradient). Secondly, we observe that the difference in convergence speeds is most pronounced when the batch size increases. Hence, being robust to outliers seems not to be solvable only by increasing the batch size and thus decreasing the noise of the mini-batch gradient. In fact, the contrary seems to be the case. This is similar to the observations made in (Kunstner et al., 2023).
Figure 7: Encoder transformer on PTB dataset, with weights fixed at initialization. We use momentum, $\mathrm{clip}_{\tau,1}$ and $\mathrm{clip}_{\tau,2}$ to estimate the full-batch gradient.
Figure 8: Encoder transformer on Wikitext-2 dataset, with weights fixed at initialization. We use momentum, $\mathrm{clip}_{\tau,1}$ and $\mathrm{clip}_{\tau,2}$ to estimate the full-batch gradient.
C.2 Learning Rate Values for Transformer Training
Here, we report for each of the language modelling tasks, the learning rate value that is displayed in Fig. 5. For each method, we tuned the learning rate over a set of values $10^{\frac{j}{2}}$ for a suitable interval of integer numbers $j\in\mathbb{Z}$.
Table 1: Tuned learning rate values used for each method. All methods use constant learning rates. For SGD-M, tuning information is also reported in Section C.1 in (Kunstner et al., 2023)).
Name SGD-M clipped-SGD VClip CClip Adam
PTB 0.1 0.316 3.16 0.316 0.001
WikiText-2 0.0316 0.316 1.0 0.316 0.001
SQuAD 0.0316 0.1 0.316 0.1 0.000316